### v3

# Max-Rank: Efficient Multiple Testing for Conformal Prediction

Abstract
Multiple hypothesis testing (MHT) commonly arises in various scientific fields, from genomics to psychology, where testing many hypotheses simultaneously increases the risk of Type-I errors. These errors can mislead scientific inquiry, rendering MHT corrections essential. In this paper, we address MHT within the context of conformal prediction, a flexible method for predictive uncertainty quantification. Some conformal prediction settings can require simultaneous testing, and positive dependencies among tests typically exist. We propose a novel correction named max-rank that leverages these dependencies, whilst ensuring that the joint Type-I error rate is efficiently controlled. Inspired by permutation-based corrections from Westfall and Young (1993), max-rank exploits rank order information to improve performance, and readily integrates with any conformal procedure. We demonstrate both its theoretical and empirical advantages over the common Bonferroni correction and its compatibility with conformal prediction, highlighting the potential to enhance predictive uncertainty quantification.
1 INTRODUCTION
Multiple hypothesis testing (MHT) occurs whenever we attempt to draw conclusions over multiple statistical hypothesis at once (Shaffer, 1995). Such scenarios frequently occur in various scientific domains, ranging from medicine and genomics (Ludbrook, 1994; Qu et al., 2010; Storey and Tibshirani, 2003) to economics and psychology (Blakesley et al., 2009; List et al., 2019). A crucial issue with MHT is that as the number of tested hypotheses increases, so does the likelihood of committing a Type-I error, i.e., erroneously rejecting a null in favour of significance. Such false positives may lead scientists astray by drawing incorrect conclusions and pursuing costly but statistically meaningless studies (Benjamini, 2010). Thus, a large body of research in statistics has preoccupied itself with multiple testing corrections to permit valid statistical inference even in such scenarios (Bender and Lange, 2001).
In this work, we investigate a particular instance of the multiple testing problem: predictive uncertainty quantification via conformal prediction (Shafer and Vovk, 2008; Vovk et al., 2005). Conformal prediction is an efficient way to produce statistically rigorous uncertainties for complex predictive models, such as deep neural networks (Angelopoulos et al., 2023). MHT issues can arise from the need to run multiple conformal procedures at once (e.g., in multi-target regression (Messoudi et al., 2021)), necessitating a correction. Moreover, these conformal procedures are known to exhibit positive dependency (Bates et al., 2022), rendering simple corrections such as Bonferroni inefficient a priori.
Exploiting the natural connections between conformal prediction and permutation testing, we propose a simple MHT correction that is able to exploit the positive dependency between hypotheses when apparent through rank ordering. Furthermore, it ensures strong Type-I error control on the family-wise error rate (FWER), i.e., the probability of a single false discovery. Our approach is closely connected to the permutation-based corrections of Westfall and Young (1993), and readily integrates with the conditions assumed by conformal prediction. To summarize, our contributions include:
  • Proposing an efficient correction called max-rank for MHT settings with positive dependency (Algo. 2) and establishing its connection to the correction procedures of Westfall and Young (1993) (§ 3.2).
  • Proving that max-rank is superior to the Bonferroni correction in terms of FWER control via multivariate copula arguments (§ 4, § 5).
  • Demonstrating both theoretically (§ 6.1) and empirically (§ 6.2) max-rank’s suitability for conformal prediction.
2 BACKGROUND & RELATED WORK
We begin by commenting on our multiple testing setting and forms of Type-I error control, with a closer look at the Bonferroni correction. We then thoroughly introduce the conformal prediction setting and its connection to hypothesis testing.
Multiple testing in parallel and in series. A useful distinction can be made between statistical testing in parallel and in series (Good, 1958). When testing in parallel, multiple hypotheses are examined simultaneously. Frequently, conclusions are drawn on the basis of the same underlying evidence such as a single common dataset (Vovk et al., 2022), making independence assumptions harder to justify. This is our setting of interest, and the one underlying conformal prediction and any arising MHT issues (see § 6). In contrast, testing in series suggests conclusions are drawn one after another, and the outcome of a test may determine whether and which subsequent hypotheses should be explored. For example, a clinical drug trial with multiple phases may only initiate a follow-up if the results are satisfactory during the first phase (Bauer, 1991).
Family-wise error rate control. We may distinguish two forms of controlling the Type-I error rate during testing: control in the strong sense via the family-wise error rate (FWER), and in the weaker sense via the false discovery rate (FDR, see § C.1). FWER control, the more stringent of the two, aims to limit the probability of committing at least one Type-I error across all considered hypotheses. If we test $m$ hypotheses and denote by $m_{\text{FP}}$ the number of falsely rejected null hypotheses (i.e., false positives), then FWER control at significance level $\alpha$ ensures that
$\mathbb{P}(m_{\text{FP}}\geq 1)\leq\alpha.$ (1)
Such strict error control is appropriate in settings where false discoveries may have strong implications, such as in drug testing and clinical trials. Since conformal prediction is oftentimes motivated from a safety perspective by its provided coverage guarantees (see § 2.1) and can be applied to similar settings (e.g., Lei and Candès (2021); Candès et al. (2023); Yin et al. (2024)), FWER control is also an appropriate notion here.
FWER control via the Bonferroni correction. Common corrections with FWER control include Bonferroni (Sedgwick, 2012), Bonferroni-Holm (Aickin and Gensler, 1996), or their Šidák-adjusted versions (Abdi et al., 2007), which all presume independence. Arguably the simplest and most widely employed, the classic Bonferroni correction adjusts each individual significance level by the (fixed) number of tested hypotheses, i.e., $\alpha_{\text{Bonf}}=\alpha/m$. Then, FWER control can be shown directly using an union bound argument (see § C.2). However, the correction is well-known to be overly conservative under positive dependency of hypotheses (Vovk et al., 2022; Benjamini, 2010), and requires specifying the number of tests $m$ beforehand.
Multiple testing under dependency. Albeit MHT issues have been addressed as early as Tukey (1953), the research on valid and efficient corrections under dependency remains remarkably active. Westfall and Young (1993) suggested an early resampling-based procedure for capturing data dependencies, while recent years have seen further work on arbitrary dependencies (Causeur et al., 2009; Sun and Tony Cai, 2009; Leek and Storey, 2008; Liu and Xie, 2019; Finner et al., 2017; Gou and Tamhane, 2018; Chi et al., 2022) and p-value merging (Wilson, 2019; Vovk et al., 2022; Chen et al., 2023). Indeed, a remark by Chi et al. (2022) that “the gap between theory and practice is mildly unsettling” suggests these problems remain far from fully solved.
Multiple testing in conformal prediction. Conformal applications necessitating some form of multiple testing correction include multi-step time series (Stankeviciute et al., 2021; Sun and Yu, 2023; Xu et al., 2024) and trajectory estimation (Lindemann et al., 2023; Muthali et al., 2023); multi-target classification (Cauchois et al., 2021; Fisch et al., 2021) and regression (Messoudi et al., 2021, 2022; Feldman et al., 2023), including conformal object detection (Andéol et al. (2023); Mukama et al. (2024), see § 6.2); treatment analysis (Lei and Candès, 2021; Yin et al., 2024) and screening (Jin and Candès, 2023; Fisch et al., 2022), or even accelerated inference (Schuster et al., 2021). Throughout, employed corrections range from Bonferroni (Stankeviciute et al., 2021; Lindemann et al., 2023; Muthali et al., 2023; Andéol et al., 2023; Jin and Candès, 2023) to Copulas (Messoudi et al., 2021; Sun and Yu, 2023), Simes (Fisch et al., 2021) and other forms of sequential testing (Lei and Candès, 2021; Yin et al., 2024; Cauchois et al., 2021; Angelopoulos et al., 2021), showcasing that a variety of solutions have been considered.
2.1 Conformal prediction
Algorithm 1 Split conformal prediction (general case)
1:Input: Pre-trained predictor $\hat{f}$, calibration data $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$, test sample $(X_{n+1},Y_{n+1})\sim\mathcal{P}_{XY}$, miscoverage rate $\alpha\in(0,1)$.
2:Output: Prediction set $\hat{C}(X_{n+1})$ for test sample.
3:Procedure:
4:Define a scoring function $s:\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ applied to $\mathcal{D}_{cal}$, resulting in a set of nonconformity scores $S=\{s(\hat{f}(X_{i}),Y_{i})\}_{i=1}^{n}=\{s_{i}\}_{i=1}^{n}$. The score $s_{i}$ encodes a notion of dissimilarity or nonconformity between the prediction $\hat{f}(X_{i})$ and true value $Y_{i}$.
[…]
6:For the new test sample $(X_{n+1},Y_{n+1})$, a valid prediction set for $X_{n+1}$ is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:\,s(\hat{f}(X_{n+1}),y)\leq\hat{Q}(1-\alpha;S)\}$. Validity refers to satisfying Eq. 2.
7:End procedure
Conformal prediction (CP) is a distribution-free, post-hoc, and finite-sample approach to predictive uncertainty quantification (Vovk et al., 2005; Shafer and Vovk, 2008). The obtained uncertainty around a predictor’s outputs is expressed in terms of prediction sets, and provides an associated guarantee on coverage (Eq. 2). The general algorithmic recipe for split CP (Papadopoulos et al., 2007), its most popular variant, is outlined in Algo. 1. Split CP requires a separate calibration dataset $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$ on which dissimilarity measures (so-called nonconformity scores) between predicted and true values are computed. For a new test sample $(X_{n+1},Y_{n+1})$, the prediction set is then constructed by including any predictions that align with the previously observed scores on $\mathcal{D}_{cal}$, as determined by a conformal quantile $\hat{Q}_{1-\alpha}$ that ensures a coverage rate of $(1-\alpha)$ for true values.
The approach is distribution-free since it does not impose any distributional assumptions. Instead, the key requirement is the exchangeability of $\mathcal{D}_{cal}\,\cup\,\{(X_{n+1},Y_{n+1})\}$, a relaxed i.i.d. assumption suggesting permutation invariance of the samples. It is post-hoc because it can be used with any ‘black-box’ pre-trained predictor without the need for training access. And it is finite-sample because the obtainable coverage guarantees do not rely on asymptotics. Specifically, if the conditions of CP are met, a probabilistic statement on containment of $Y_{n+1}$ in the prediction set $\hat{C}(X_{n+1})$ for some tolerated miscoverage rate $\alpha\in(0,1)$ is given as
$\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1}))\geq 1-\alpha.$ (2)
Importantly, the provided coverage guarantee is marginally valid, i.e., it only holds on average across test and calibration data and may thus fail to cover for any particular observation.
Conformal prediction as hypothesis testing. The CP procedure in Algo. 1 can also be interpreted from a hypothesis testing perspective (Shi et al., 2013; Bates et al., 2022; Vovk et al., 2005). Given a prediction set candidate value $y\in\mathcal{Y}$, a statistical hypothesis for the test sample is established as $H_{0}:Y_{n+1}=y,\,\,H_{1}:Y_{n+1}\neq y$, with test statistic $s_{n+1}=s(\hat{f}(X_{n+1}),y)$ equal to its nonconformity score. The set of calibration scores $S=\{s_{1},\dots,s_{n}\}$ provides an empirical null distribution to compute a p-value $\hat{P}_{n+1}(y;S)$ as
$\hat{P}_{n+1}(y;S)=|\{i=1,\dots,n:s_{i}\geq s_{n+1}\}|/(n+1),$ (3)
i.e., the fraction of samples which conform worse than $s_{n+1}$. A testing decision on not rejecting $H_{0}$ – that is, inclusion in the prediction set – is taken at significance level $\alpha$ as $\hat{P}_{n+1}(y;S)>\alpha$. This approach can be interpreted as permutation testing: we are assessing if the observed nonconformity for our candidate value does not stray too much from previous observations, and the sample $(X_{n+1},y)$ thus permutes with $D_{cal}$. We may then interpret running multiple conformal procedures at once as permutation testing in parallel, typically at identical significance level $\alpha$ across tests.
Equivalence of events. By the prediction set construction outlined in Algo. 1 and the above testing perspective, a useful equivalence of events for some set candidate value $y\in\mathcal{Y}$ can be established as
$y\in\hat{C}(X_{n+1})$ $\Leftrightarrow s_{n+1}\leq\hat{Q}(1-\alpha;S)\Leftrightarrow\hat{P}_{n+1}(y;S)>\alpha,\text{}$
$y\notin\hat{C}(X_{n+1})$ $\Leftrightarrow s_{n+1}>\hat{Q}(1-\alpha;S)\Leftrightarrow\hat{P}_{n+1}(y;S)\leq\alpha.$ (4)
Notably, the conformal quantile $\hat{Q}_{1-\alpha}$ [^1] We use $\hat{Q}_{1-\alpha}$ as a shortened notation for $\hat{Q}(1-\alpha;S)$. ensuring a fraction of $(1-\alpha)$ values below it is the equivalent of a permutation test at significance level $\alpha$. In both cases, the Type-I error is controlled at the same level $\alpha$. We further comment on the quantile’s properties in the following remarks:
Remark 1.
The Type-I error control property of $\hat{Q}_{1-\alpha}$ at level $\alpha$ can be expressed in terms of coverage, i.e., it is the quantile which ensures that at least a fraction $(1-\alpha)$ of values are below and thus covered by it.
Remark 2.
The smallest possible quantile $\hat{Q}_{1-\alpha}$ which maintains Type-I error control at level $\alpha$ is tight in terms of coverage, i.e., it provides the smallest and thus most efficient conformal prediction sets.
3 MULTIPLE TESTING CORRECTION VIA MAX-RANK
We next formalize and describe our proposed multiple testing correction max-rank, and thoroughly connect it to corrections by Westfall and Young (1993) in § 3.2.
Notation. For consistency, we denote matrices as upper-case bolded ($\mathbf{M}$), vectors as lower-case bolded ($\mathbf{v}$), scalars without bolding ($s$). A multivariate random variable may also be upper-case bolded given the context ($\mathbf{X}$). An index set $\{1,\dots,n\}$ is abbreviated $[n]$ for any $n\geq 1$. Similarly, a sequence $x_{1},\dots,x_{n}$ is shortened to $x_{1:n}$. The rank of an element $x_{i}$ in a set $\{x_{1:n}\}$ of $n$ distinct elements is given by
[…]
i.e., the size of the set of elements below $x_{i}$. Finally, the $l^{\infty}$-norm is given by $\lVert\mathbf{x}\rVert_{\infty}=\max_{1\leq i\leq n}|x_{i}|$ for a discrete vector $\mathbf{x}$ of size $n$.
3.1 The max-rank algorithm
Consider running $m$ hypothesis tests in parallel, and for any particular testing dimension $k\in[m]$ the statistics forming an empirical null distribution are given by $\textbf{s}_{k}\in\mathbb{R}^{n}$. In the conformal setting, $\mathbf{s}_{k}$ equates the set of nonconformity scores on $\mathcal{D}_{cal}$. We now desire finding adjusted significance levels $\alpha_{k}$ for all dimensions such that the FWER is globally controlled at target level $\alpha$.
We algorithmically describe our proposed max-rank correction in Algo. 2. First, the empirical nulls of equal size are collected in a matrix of test statistics $\mathbf{S}=[s_{i,k}]\in\mathbb{R}^{n\times m}$, where $i\in[n]$ and $k\in[m]$. The $m$ columns span the tested hypotheses, and the $n$ rows represent values in each null distribution. Next, we obtain a similar matrix containing the column-wise ranks of $\mathbf{S}$ as $\mathbf{R}=[r_{i,k}]\in\mathbb{N}^{n\times m}$, where also $i\in[n]$ and $k\in[m]$. By Eq. 4, the search for a significance level $\alpha_{k}$ for each test is equivalent to computing the conformal quantile $\hat{Q}(1-\alpha;\mathbf{s}_{k})$ in the space of test statistics. This in turn equates determining the value $s_{t,k}$, where $t=\hat{Q}(1-\alpha;\mathbf{r}_{k})$ denotes a quantile in rank space ensuring the desired error control (Remark 1). max-rank adjusts these rank quantiles for FWER control by forming a composite empirical null $\mathbf{r}_{\max}$ via the $l^{\infty}$-norm across testing dimensions, and using the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ to inform globally corrected significance levels for each hypothesis test $k\in[m]$.
Algorithm 2 Multiple testing correction via max-rank
1:Input: Hypothesis tests $k=1,\dots,m$ and associated empirical nulls $\mathbf{s}_{1},\dots,\mathbf{s}_{m}$ of equal size $n$.
2:Output: Adjusted quantiles with FWER control at level $\alpha$ as $(\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m})\in\mathbb{R}^{m}$.
3:Procedure:
4:Collect empirical nulls in a matrix of test statistics $\mathbf{S}\in\mathbb{R}^{n\times m}$, and their respective column-wise ranks in a matrix of rank statistics $\mathbf{R}\in\mathbb{N}^{n\times m}$.
5:Apply the $l^{\infty}$-norm row-wise to the ranks. Let $\mathbf{r}_{i}=(r_{i,1}\cdots r_{i,m})$ be the $i$-th row of $\mathbf{R}$, then $\mathbf{r}_{\max}=(\lVert\mathbf{r}_{1}\rVert_{\infty},\dots,\lVert\mathbf{r}_{n}\rVert_{\infty})^{T}=(\underset{1\leq k\leq m}{\max}|r_{1,k}|,\cdots,\underset{1\leq k\leq m}{\max}|r_{n,k}|)^{T}\in\mathbb{N}^{n}.$
6:Compute the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ as the $\lceil(n+1)(1-\alpha)/n\rceil$-th empirical quantile of $\mathbf{r}_{\max}$.
7:Sort $\mathbf{S}$ column-wise ascending such that $s_{1,k}<\dots<s_{n,k}$ for every $k\in[m]$.
8:Select the test statistic $s_{r_{\max},k}$ at rank $r_{\max}$ column-wise in $\mathbf{S}$ as the adjusted quantile $\hat{Q}_{1-\alpha,k}$ for every test dimension $k\in[m]$.
9:End procedure
Intuition. Our max-rank correction intuitively collapses multiple empirical nulls into a single ‘worst-case’ null for which the error-controlling quantile is computed. Thus, any potential dependencies across test statistics which are reflected in the rank ordering are efficiently aggregated in the ‘max-rank’ statistic $\mathbf{r}_{\max}$, and subsequently the selected rank quantile $r_{\max}$. By virtue of applying the $l^{\infty}$-norm, $r_{\max}$ ensures that adjusted significance levels provide FWER control.
If the empirical nulls are uncorrelated and the tested hypotheses can thus be considered independent, no benefit is obtained from max-rank’s aggregation step. We prove that in such cases our approach performs at least on par with the Bonferroni correction (§ 4), and never worse. On the other hand, if the tests are perfectly correlated each empirical null exhibits identical rank ordering, and max-rank will provide the optimal quantile solution for each individual testing dimension. For any cases with positively dependent tests inbetween the two extremes, our approach ensures superior performance over Bonferroni, and scales well even as the number of hypotheses grows (see § 5).
3.2 Relation to Westfall & Young corrections
A strong connection is apparent between max-rank and the max-T and min-P resampling-based corrections of Westfall and Young (1993), which offer similar FWER control. The max-T method provides adjusted p-values by comparing test statistics to the distribution of maximum statistics across hypotheses (Algo. 3), while the min-P method adjusts p-values based on the minimum observed p-values after resampling (Algo. 4). In both cases, existing dependency structure between tests is leveraged through repeated data resampling.
Exploring these two corrections through the lens of our conformal prediction setting, we find that max-rank can be cast as a procedure that (1) naturally extends max-T to the rank space, and (2) describes a particular instance of min-P arising from our problem setting. As such, we argue that max-rank exhibits the advantages of both methods by (1) permitting its direct application to obtained test statistics as in max-T, whilst (2) remaining scale-invariant as in min-P. We next relate to each correction in more detail.
Relation to max-T. Our conformal setting directly provides empirical null distributions for every hypothesis by use of the calibration set. The computed nonconformity scores on $\mathcal{D}_{cal}$ leverage available labels to produce statistics satisfying the tested null $H_{0}:Y_{n+1}=y$, and are thus interpretable as an analog to the required “data resampling under the null” (Westfall and Young, 1993). It follows that the adjusted p-values in Algo. 3 match the conformal p-values in Eq. 3 simply by replacing the score set $S$ with a maximum score set $S_{\max}$, i.e., replacing $\hat{P}_{n+1}(y;S)$ with $\hat{P}_{n+1}(y;S_{\max})$ [^2] And correcting for $(n+1)$ samples instead of $n$.. Instead of adjusting the p-values as in Algo. 3 we can directly adjust the significance level $\alpha$, which by Eq. 4 translates to computing the quantile $\hat{Q}(1-\alpha;S_{\max})$ instead of $\hat{Q}(1-\alpha;S)$. Comparing this to max-rank, we notice that the difference lies in its direct application to the maximum statistics, rather than translating them into rank space first (i.e., using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$). This means that max-T is sensitive to scale and requires test statistics across hypotheses to be directly comparable. Interestingly, such a “maximum of scores” approach has been intuitively employed by researchers in conformal prediction, e.g., Andéol et al. (2023); Schuster et al. (2021); Fisch et al. (2022), but never been recognized as an application case of Westfall and Young (1993).
Relation to min-P. By the same resampling interpretation as above, unadjusted p-values for each hypothesis in Algo. 4 can be computed via conformal p-values in Eq. 3. Using the rank definition in Eq. 5, a given p-value $\hat{P}_{n+1}(y;S)$ can be expressed in terms of score ranking as $((n+1)-\text{rank}(s_{n+1};S))/(n+1)$ [^3] Observe that the permutation-based p-value in Eq. 3 counts the fraction of values above the quantity of interest, whereas the rank counts values below it.. Thus, computing the minimum p-values for min-P can be interpreted as taking the maximum over inverted and normalized ranks. Using Eq. 4, we again translate adjusting the p-values as adjusting the significance level via $\hat{Q}(\alpha;\mathbf{p}_{\min})$ instead. Note that this quantile is now taken over (min.) p-values, and thus only permits testing in p-value space. In contrast, max-rank maintains the simplicity of operating directly on the statistics, facilitating the interpretation of testing decisions.
4 THEORETICAL RESULTS
We next briefly describe our theoretical results for max-rank, deferring details and proofs to Appendix A. The multiple testing problem is first rephrased in terms of multivariate copulas, and our two key results state that (1) max-rank provides a valid solution with FWER control, and (2) is assured to perform better than Bonferroni under positive dependency of tests.
Recasting the multiple testing problem (§ A.1). Consider an $m$-dimensional random variable $\mathbf{X}\in\mathbb{R}^{m}$ and denote its c.d.f. as $F_{\mathbf{X}}(\mathbf{x})$. Using the interpretation in Remark 1, determining adjusted significance levels $(\alpha_{1},\dots,\alpha_{m})$ with joint Type-I error control at level $\alpha$ is then the equivalent to finding a vector $\mathbf{x}_{*}\in\mathbb{R}^{m}$ satisfying $F_{\mathbf{X}}(\mathbf{x}_{*})\geq 1-\alpha$. By using Sklar’s Theorem (Sklar, 1973) to express the c.d.f. in terms of a copula $C_{\mathbf{X}}$, and defining a vector of probabilities $(q_{*,1},\dots,q_{*,m})=(F_{X_{1}}(x_{*,1}),\dots,F_{X_{m}}(x_{*,m}))\in[0,1]^{m}$ restricted to a fixed value $q_{*}=q_{*,1}=\dots=q_{*,m}$, we can restate the problem as the search for a solution
$\min q_{*}\text{ subject to }C_{\mathbf{X}}(q_{*},\dots,q_{*})\geq 1-\alpha,$ (6)
where the $\min$ is motivated by our efficiency goals to best match target coverage $(1-\alpha)$ (Remark 2).
Obtaining a solution via max-rank (§ A.2). For a given solution to Eq. 6 the desired quantiles with FWER control are then computed as $\mathbf{x}_{*}=(F_{X_{1}}^{-1}(q_{*}),\dots,F_{X_{m}}^{-1}(q_{*}))\in\mathbb{R}^{m}$. These are equal to the quantiles $\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m}$ described in Algo. 2. Leveraging operations in the rank space, we show that max-rank provides such a solution $q_{\max}$. We state this in the following proposition:
Proposition 1.
The max-rank procedure (Algo. 2) provides a solution $q_{\max}$ to the constrained problem in Eq. 6 with FWER control at level $\alpha$.
Furthermore, the $\min$ implies that the associated quantiles $\mathbf{x}_{\max}$ are optimally tight in terms of coverage, under the condition that $q_{\max}=q_{\max,1}=\dots=q_{\max,m}$.
Comparison to the Bonferroni correction (§ A.3). Under independence the c.d.f. $F_{\mathbf{X}}(\mathbf{x}_{*})$ factorizes, and the coverage constraint in Eq. 6 simplifies to ensuring $(q_{*})^{m}\geq 1-\alpha$. max-rank satisfies the condition directly by Prop. 1, i.e., we have $q^{\perp}_{\max}\geq(1-\alpha)^{1/m}$ denoting the solution under independence. In contrast, Bonferroni ensures the condition is met by $q_{\text{Bonf}}=(1-\alpha/m)$, since $(1-\alpha/m)\geq(1-\alpha)^{1/m}$ for any $m\geq 1,\,\alpha\in(0,1)$. It directly follows that $q^{\perp}_{\max}\leq q_{\text{Bonf}}$. In the case of positive dependency we impose that correlated tests follow positive lower orthant dependency (Nelsen, 2006), suggesting a higher probability of taking on similar values than under independence. This directly implies that $q_{\max}\leq q^{\perp}_{\max}$. We summarize these results as:
Proposition 2.
The solution $q_{\max}$ provided by the max-rank procedure (Algo. 2) to Eq. 6 ensures that we have $q_{\max}\leq q^{\perp}_{\max}\leq q_{\text{Bonf}}$ in the case of both independent and positively dependent test statistics.
[…]
5 SIMULATION STUDY
[…]
Global FWER control. We compare the adjusted significance levels produced by both max-rank and Bonferroni across varying correlation strengths and for $m\in\{1,3,5,10,20\}$ in Fig. 1. We observe two key drawbacks for Bonferroni. Firstly, the correction becomes increasingly conservative as correlation between tests increases, with the adjusted level $\hat{\alpha}$ tending towards zero. Secondly, the rate accelerates as the number of tests increases. In contrast, max-rank suffers from neither of these pathologic behaviours and is able to tightly maintain FWER control at target level $\alpha$, ensuring good performance independently of both correlation strength and test dimensionality (Prop. 1).
Individual Type-I error control. We next compare the adjusted significance levels obtained by max-rank and Bonferroni for individual testing dimensions $k\in[m]$ and for fixed $m=5$ against varying correlation strengths in Fig. 2. Bonferroni limits individual Type-I error levels to $\hat{\alpha}_{\text{Bonf}}=\alpha/m=0.01$ in order to achieve FWER control, forcing overly conservative behaviour in each test. In contrast, max-rank is able to exploit any positive dependency between tests, resulting in reduced requirements on individual Type-I error levels as the correlation increases (Prop. 2). This, of course, strongly benefits the chances of a significant result in each dimension. Meanwhile, we observe in Fig. 1 for $m=5$ that the FWER remains controlled.
[…]
6 APPLICATION TO CONFORMAL PREDICTION
We now elaborate on how our proposed max-rank correction is useful for the conformal prediction setting introduced in § 2.1. First, we illustrate how multiple testing issues arise whenever multiple CP procedures are run in parallel. We then theoretically motivate why max-rank integrates well with the conformal setting, and experimentally validate its use in two conformal applications requiring such a correction.
Multiple testing issues in conformal prediction. We previously motivated that running multiple conformal procedures at once equates permutation testing in parallel at significance level $\alpha$ (Eq. 3). By Eq. 4 and Remark 1 this is the same as ensuring $(1-\alpha)$ coverage for an $m$-dimensional response $\mathbf{Y}_{n+1}=(Y_{n+1,1},\dots,Y_{n+1,m})$ simulatenously. Take $\mathbf{X}_{n+1}=(X_{n+1,1},\dots,X_{n+1,m})$ to be a matching multivariate input, then we directly observe how the coverage guarantee in Eq. 2 is violated since
$\mathbb{P}\left(\mathbf{Y}_{n+1}\in\hat{C}(\mathbf{X}_{n+1})\right)$ $=\mathbb{P}\left(\bigcap_{k=1}^{m}\left(Y_{n+1,k}\in\hat{C}(X_{n+1,k})\right)\right)$
$\geq 1-\mathbb{P}\left(\bigcup_{k=1}^{m}\left(Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\right)\right)$
$\geq 1-\sum_{k=1}^{m}\mathbb{P}\left(Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\right)$
$=1-\sum_{k=1}^{m}\mathbb{P}\left(\hat{P}_{n+1,k}(Y_{n+1,k};S)\leq\alpha_{k}\right)$
$\geq 1-\sum_{k=1}^{m}\alpha_{k}=1-m\alpha\quad\text{for }\alpha_{k}=\alpha\text{ fixed.}$. (7)
That is, guaranteed coverage is limited to $1-m\alpha\leq 1-\alpha$ for any number of tests $m\in\mathbb{N}^{+}$, requiring a correction.
A key insight is that due to their joint dependence on common calibration data $\mathcal{D}_{cal}$, any obtained conformal p-values are bound to exhibit some degree of positive dependency. This behaviour is referred to as positive regression dependent on a subset (PRDS) (Bates et al., 2022; Benjamini and Yekutieli, 2001; Chi et al., 2022), and highlights that any multiple testing correction relying on an independence assumption (such as Bonferroni) is bound to be overly conservative a priori.
6.1 Using max-rank for conformal prediction
Since we have intimately connected max-rank’s formulation to conformal prediction thus far, we may rightly assume that it integrates well with any CP algorithm following Algo. 1. In order to rigorously verify this fact, we assert that max-rank satisfies two key requirements imposed by conformal prediction: (1) the exchangeability condition on the underlying data is preserved, and (2) Eq. 7 is corrected in such a way that the conformal coverage guarantee from Eq. 2 is upheld.
We demonstrate in § A.4 that max-rank satisfies both conditions by looking at its two key operations, namely (1) operating in the rank space of nonconformity scores, and (2) applying a $\max$-operation to these ranks (or more generally, any set of exchangeable random variables). While the requirement on valid coverage directly follows from max-rank’s FWER-controlling property (Prop. 1), an additional proof is given via rank coverage. We summarize these insights as follows:
Proposition 3.
The max-rank procedure (Algo. 2) provides a multiple testing correction for conformal prediction that preserves exchangeability and ensures valid coverage (Eq. 2).
6.2 Conformal prediction experiments
Finally, we showcase max-rank’s usefulness experimentally for two applications, conformal multi-target regression and conformal object detection.
Experiment design. We consider applying conformal prediction to multi-target regression problems where $m$ real-valued and possibly dependent targets are predicted simultaneously. We aim to ensure that all $m$ targets are jointly covered with probability $(1-\alpha)$ by our conformal procedure, and consider a range of multiple testing corrections for that purpose: Bonferroni and its slightly more powerful Šidák-adjustment (Abdi et al., 2007), two copula-based corrections from Messoudi et al. (2021), max-T (Westfall and Young, 1993) and finally max-rank. We include the copula-based corrections because they also address settings of positive dependency, and since max-rank’s theoretical motivation via copulas provides an interesting link. For CP (Algo. 1) we employ simple nonconformity scores of the form $|Y_{i,k}-\hat{f}(X_{i,k})|$ where $i\in[n],k\in[m]$, and construct the prediction intervals for test points as $\hat{C}(X_{n+1,k})=[\hat{f}(X_{n+1,k})-\hat{Q}_{1-\hat{\alpha},k},\,\hat{f}(X_{n+1,k})+\hat{Q}_{1-\hat{\alpha},k}]$, where $\hat{Q}_{1-\hat{\alpha},k}$ now denotes the conformal quantile at a possibly adjusted individual significance level $\hat{\alpha}_{k}$. We compare results with different corrections across two key conformal metrics (Vovk et al., 2005): empirical coverage as a measure of validity to verify Eq. 2, and average prediction interval width to assess the efficiency of the solution. See also Appendix D.
Figure 3: Empirical coverage (left) and mean prediction interval width (right) for different corrections ($\alpha=0.1$) on BDD100k. Results are averaged across objects from multiple classes (see § D.5), while boxplots depict the distribution over $100$ random trials. Inference times in sec. per trial (left to right by method): $2.7$, $2.7$, $17.8$, $62.6$, $2.7$, and $2.95$. max-rank outperforms other corrections with notably lower runtimes than copula-based methods.
scpf ($m=3$) rf1 ($m=8$) scm1d ($m=16$)
Correction Coverage Interval Width Coverage Interval Width Coverage Interval Width
No correction $0.83\pm 0.04$ $18.89\pm 3.62$ $0.50\pm 0.01$ $1.88\pm 0.05$ $0.50\pm 0.02$ $243.87\pm 4.50$
Bonferroni $0.95\pm 0.02$ $46.73\pm 13.39$ $0.92\pm 0.01$ $7.18\pm 0.44$ $0.96\pm 0.01$ $793.70\pm 50.65$
Bonf.-Šidák $0.95\pm 0.02$ $46.73\pm 13.39$ $0.92\pm 0.01$ $6.98\pm 0.43$ $0.96\pm 0.01$ $793.70\pm 50.65$
Gumbel Copula $0.90\pm 0.03$ $\bm{30.15\pm 5.53}$ $0.91\pm 0.01$ $6.55\pm 0.39$ $0.94\pm 0.01$ $697.89\pm 36.48$
Emp. Copula $0.92\pm 0.03$ $34.38\pm 6.74$ $0.91\pm 0.01$ $\underline{6.32\pm 0.31}$ $0.91\pm 0.01$ $\underline{579.21\pm 15.43}$
max-T $0.91\pm 0.03$ $53.77\pm 10.63$ $0.90\pm 0.01$ $6.76\pm 0.24$ $0.90\pm 0.01$ $581.93\pm 14.14$
max-rank (Ours) $0.91\pm 0.03$ $\underline{32.28\pm 6.26}$ $0.90\pm 0.01$ $\bm{6.08\pm 0.29}$ $0.90\pm 0.01$ $\bm{564.01\pm 13.79}$
Table 1: Empirical coverage and mean prediction interval width for different corrections ($\alpha=0.1$) on three multi-target datasets: scpf (1137 samples), rf1 (9125 samples), and scm1d (9803 samples). We employ a random forest regressor (Pedregosa et al., 2011) and $20\%$ calibration split, and report means and std. deviation across $100$ random trials. max-rank consistently provides strong results (best results, second best).
Conformal multi-target regression. Following Messoudi et al. (2021) we consider three tabular datasets with a varying number of targets [^4] Obtainable at https://github.com/tsoumakas/mulan: scpf (website traffic data with $m=3$), rf1 (river flow data with $m=8$), and scm1d (product pricing data with $m=16$). Our results for the usual target coverage of $(1-\alpha)=90\%$ are displayed in Table 1. Firstly, we observe that all corrections provide valid coverage, whereas without we (strongly) undercover. Among corrections, Bonferroni and Bonferroni-Šidák provide overly conservative results, and copula methods perform generally well. In particular for the scpf dataset the parametric Gumbel copula benefits from the low test dimensionality ($m=3$), while max-T and max-rank suffer from its small sample size. Overall, max-rank tends to outperform other corrections in terms of interval width, and matches the target coverage level very closely.
Conformal object detection. We also explore the more complex task of predicting an object’s location in an image via its bounding box coordinates ($m=4$) with a deep learning-based object detector (Wu et al., 2019). Following Mukama et al. (2024) we use the large-scale BDD100k dataset (Yu et al., 2020), but apply CP separately to multiple classes (as opposed to only considering a single class ‘car’) [^5] In concurrent work, we develop a richer conformal methodology for uncertainty in object detection and use max-rank as a correction (Timans et al., 2024).. Our results in Fig. 3 stress that max-rank ensures both tight empirical coverage and the smallest prediction intervals across considered methods. Copula-based corrections provide competitive results but suffer from substantially larger computational requirements. The Gumbel copula requires numerical optimization to determine its parametric form, whereas the empirical copula repeatedly evaluates ranks across all $m$ dimensions during inference. In contrast, max-rank collapses the test dimensions and its simple quantile operations result in low runtimes comparable to Bonferroni and max-T.
7 DISCUSSION
We introduced max-rank, an efficient MHT correction for testing under positive dependency which tailors Westfall and Young (1993)’s corrections for conformal prediction. While we provide strong theoretical and empirical results, our method’s practicality is restricted to (1) data dependencies that become apparent through rank ordering, (2) settings of positive dependency, since we do not benefit from rank information under negative dependency and thus do not improve upon Bonferroni, and (3) its permutation-based, data-driven nature necessitating a sufficient ratio of test statistics $n$ to test dimensions $m$ for empirical FWER control. Future work includes establishing stronger theoretical results, such as on convergence rates to target level $\alpha$, the dependence between quantities $m$, $n$, and $\alpha$, and its relationship to the empirical copula. Of course, we may also experimentally extend to further applications such as multi-step time series (Sun and Yu, 2023). Our work contributes to the ongoing research on multiple testing under dependency, all the while addressing existing needs in predictive settings using conformal prediction.
Acknowledgements
We thank Aaditya Ramdas and Jelle Goeman for making us aware of connections to the Westfall & Young testing procedures. This project was generously supported by the Bosch Center for Artificial Intelligence. Eric Nalisnick did not use any resources from Johns Hopkins for this work. Author contributions are as follows: Alexander Timans established the theoretical connections to conformal prediction and Westfall & Young, performed experiments and paper writing. Christoph-Nikolas Straehle developed the max-rank algorithm and proved several theoretical properties. Kaspar Sakmann, Christian Naesseth and Eric Nalisnick provided project guidance as well as technical and writing feedback.
Appendix A MATHEMATICAL DETAILS
This section includes the details of our theoretical results for max-rank (Prop. 1 and Prop. 2 in § 4), as well as its use for conformal prediction (Prop. 3 in § 6.1).
A.1 Recasting the multiple testing problem
We begin by reformulating our multiple testing problem in terms of multivariate copulas. We show that the search for an adjusted global significance level with FWER control can be recast as the search for a probability vector satisfying a coverage constraint under a multivariate copula.
Let $\mathbf{X}=(X_{1},\dots,X_{m})\in\mathbb{R}^{m}$ be an $m$-dimensional random variable where each dimension $k\in[m]$ corresponds to a hypothesis test. We denote the multivariate cumulative distribution function (c.d.f.) of $\mathbf{X}$ as $F_{\mathbf{X}}(\mathbf{x})=\mathbb{P}(X_{1}\leq x_{1},\dots,X_{m}\leq x_{m})$, for any realization $\mathbf{x}=(x_{1},\dots,x_{m})$. As such $F_{\mathbf{X}}(\mathbf{x})$ is right-continuous and non-decreasing in its arguments. Looking at the FWER control definition in Eq. 1 we see that $\mathbb{P}(m_{\text{FP}}\geq 1)=1-\mathbb{P}(m_{\text{FP}}=0)\leq\alpha\Leftrightarrow\mathbb{P}(m_{\text{FP}}=0)\geq 1-\alpha$, i.e., the probability of not producing any false positives or Type-I errors across test dimensions is kept above $(1-\alpha)$. Since by Remark 1 such a Type-I error control in any single test dimension can be expressed in terms of a quantile $x_{k}$ ensuring $(1-\alpha)$ coverage, a joint Type-I error control across tests matching the FWER is ensured for any quantile vector $\mathbf{x_{*}}=(x_{*,1},\dots,x_{*,m})$ satisfying
$F_{\mathbf{X}}(\mathbf{x}_{*})\geq 1-\alpha.$ (8)
Such an instance $\mathbf{x_{*}}$ always exists since $F_{\mathbf{X}}(\mathbf{x_{*}})\rightarrow 0$ if any $x_{*,k}\rightarrow-\infty$ and $F_{\mathbf{X}}(\mathbf{x_{*}})\rightarrow 1$ as all $x_{*,k}\rightarrow\infty$.
Alternatively, we can use Sklar’s Theorem (Sklar, 1973) to express the c.d.f. $F_{\mathbf{X}}(\mathbf{x})$ in terms of a multivariate copula $C_{\mathbf{X}}(\mathbf{u})$ with uniform marginals $\mathbf{u}\in[0,1]^{m}$ capturing the dependency structure among tests. That is, we can express $F_{\mathbf{X}}(\mathbf{x_{*}})$ as $C_{\mathbf{X}}(F_{X_{1}}(x_{*,1}),\dots,F_{X_{m}}(x_{*,m}))$, and if we explicitly define a vector of probabilities $\mathbf{q_{*}}=(q_{*,1},\dots,q_{*,m})=(F_{X_{1}}(x_{*,1}),\dots,F_{X_{m}}(x_{*,m}))\in[0,1]^{m}$, as $C_{\mathbf{X}}(q_{*,1},\dots,q_{*,m})$. Thus the coverage target from Eq. 8 can be re-stated as
$F_{\mathbf{X}}(\mathbf{x_{*}})=\mathbb{P}(X_{1}\leq F_{X_{1}}^{-1}(q_{*,1}),\dots,X_{m}\leq F_{X_{m}}^{-1}(q_{*,m}))=C_{\mathbf{X}}(q_{*,1},\dots,q_{*,m})\geq 1-\alpha.$ (9)
Since max-rank computes a single rank quantile which informs the conformal quantile in each test dimension, we restrict $\mathbf{q}_{*}$ to a single fixed quantile probability $q_{*}=q_{*,1}=\dots=q_{*,m}$, i.e., $\mathbf{q_{*}}=(q_{*},\dots,q_{*})$. The same restriction is made independently in Messoudi et al. (2021) to simplify copula calculations. Motivated by our efficiency goals to match target coverage $(1-\alpha)$ as closely as possible and obtain the smallest conformal prediction sets (Remark 2), we seek the smallest fixed probability $q_{*}$ satisfying Eq. 9. Thus a solution to our initial FWER control problem can be restated as the search for
$\min q_{*}\text{ subject to }C_{\mathbf{X}}(q_{*},\dots,q_{*})\geq 1-\alpha.$ (10)
Such an instance $\mathbf{q_{*}}=(q_{*},\dots,q_{*})$ also always exists since $C_{\mathbf{X}}(\mathbf{q_{*}})\rightarrow 0$ as $q_{*}\rightarrow-\infty$ and $C_{\mathbf{X}}(\mathbf{q_{*}})\rightarrow 1$ as $q_{*}\rightarrow\infty$.
A.2 Obtaining a solution via max-rank
For a given solution $q_{*}$ satisfying Eq. 10, the desired quantile vector ensuring FWER control is then computed by evaluating the probability under each inverse marginal as $\mathbf{x_{*}}=(F_{X_{1}}^{-1}(q_{*}),\dots,F_{X_{m}}^{-1}(q_{*}))\in\mathbb{R}^{m}$. Note that now the elements of $\mathbf{x_{*}}$ may differ from each other, since the marginals are not necessarily identical. The obtained quantiles in $\mathbf{x_{*}}$ are equal to the adjusted quantiles $\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m}$ described in Algo. 2. By relating the element-wise ranks of $x_{*,k}$ under $\mathbf{S}$ to the marginals $F_{X_{k}}$ and establishing a correspondence between the copula and a joint c.d.f. in rank space, we next show that our max-rank correction provides a solution $q_{\max}$ to Eq. 10.
For that, we relate the algorithmic steps in Algo. 2 to our problem setting in § A.1. Consider the matrix of test statistics $\mathbf{S}\in\mathbb{R}^{n\times m}$ first introduced in § 3, and denote by $S_{k}$ its $k$-th column corresponding to the $k$-th empirical null. We then associate $S_{k}$ with the $k$-th dimension of $\mathbf{X}$, i.e., the elements $s_{i,k}\in S_{k},\,i\in[n]$ match to $n$ realizations $\{x_{1,k},\dots,x_{n,k}\}$ of $X_{k}$, and $S_{k}$ thus provides an empirical estimate of the marginal $F_{X_{k}}$. Next, we evaluate the element-wise ranks of $\mathbf{x_{*}}$ under $\mathbf{S}$. For any element $x_{*,k}\in\mathbf{x_{*}}$ a direct correspondence between its rank under $S_{k}$ (see Eq. 5) and marginal $F_{X_{k}}$ is given by
$\text{rank}(x_{*,k};S_{k})=r_{*,k}=n\cdot q_{*}=n\cdot F_{X_{k}}(x_{*,k}),$ (11)
that is, the rank of $x_{*,k}$ is precisely the rank that ensures $q_{*}$ probability mass below it. In particular, we then have that $q_{*}=r_{*,k}/n=\tilde{r}_{*,k}$ for any $k\in[m]$, where $\tilde{r}_{*,k}\in[0,1]$ defines the normalized rank amenable to a copula. Since we have distinct scores (e.g., by introducing jitter noise) and thus distinct ranks $\{1,\dots,n\}$, it follows that the same rank $\tilde{r}_{*}=\tilde{r}_{*,1}=\dots=\tilde{r}_{*,m}$ is retrieved in every dimension. Thus, we can establish an equivalence between the copula $C_{\mathbf{X}}$ which describes the joint c.d.f. of $\mathbf{X}$, and the matching joint c.d.f. in the (normalized) rank space of $\mathbf{X}$, which we denote $F_{\tilde{\mathbf{R}}}$. We use $\tilde{\mathbf{R}}$ to stress the association with the matrix of rank statistics $\mathbf{R}\in\mathbb{N}^{n\times m}$ described in § 3. Leveraging a $\max$-operation, we can then state that
$C_{\mathbf{X}}(q_{*},\dots,q_{*})=F_{\tilde{\mathbf{R}}}(\tilde{r}_{*},\dots,\tilde{r}_{*})=\mathbb{P}(\tilde{R}_{1}\leq\tilde{r}_{*},\dots,\tilde{R}_{m}\leq\tilde{r}_{*})=\mathbb{P}(\max_{1\leq k\leq m}\tilde{R}_{k}\leq\tilde{r}_{*})=F_{\max\mathbf{\tilde{R}}}(\tilde{r}_{*}),$ (12)
where $F_{\max\mathbf{\tilde{R}}}$ now denotes the one-dimensional c.d.f. in the space of maximum ranks across dimensions $k\in[m]$. Intuitively, ensuring that the condition holds jointly for ranks in all dimensions is equivalent to satisfying it in the ‘worst’ one. $F_{\max\mathbf{\tilde{R}}}$ is precisely the space in which max-rank operates after the $l^{\infty}$-norm is applied, and in which following Algo. 2 a rank $\tilde{r}_{\max}=r_{\max}/n=q_{\max}$ with probability mass $(1-\alpha)$ below it is computed. Thus we can employ Eq. 12 in its reverse direction to claim
$1-\alpha\leq F_{\max\mathbf{\tilde{R}}}(\tilde{r}_{\max})=F_{\mathbf{\tilde{R}}}(\tilde{r}_{\max},\dots,\tilde{r}_{\max})=C_{\mathbf{X}}(q_{\max},\dots,q_{\max}),$ (13)
therefore providing a solution which satisfies the constraint in Eq. 10. Furthermore, since we precisely compute the rank at target level $(1-\alpha)$ (and ignoring any finite-sample correction), our solution $q_{\max}$ also satisfies the $\min$ in Eq. 10, and the obtained quantiles $\mathbf{x}_{\max}$ are optimally tight in terms of coverage under the imposed restriction that $q_{\max}=q_{\max,1}=\dots=q_{\max,m}$. Note that if the restriction is lifted, marginally tighter individual quantiles $x_{\max,k}$ are potentially obtainable in some dimensions. However, it becomes unclear how the required algorithmic modifications to max-rank would continue to uphold FWER control.
We summarize our arguments by restating our proposition from § 4, which follows from the above:
Proposition 1.
The max-rank procedure (Algo. 2) provides a solution $q_{\max}$ to the constrained problem in Eq. 6 with FWER control at level $\alpha$.
A.3 Comparison to the Bonferroni correction
Finally we compare our max-rank procedure to the Bonferroni correction, and demonstrate that max-rank performs either on par or better than Bonferroni under settings of independence or positive dependency between hypothesis tests. We first consider the case of independence, which the Bonferroni correction generally assumes.
Under independence. If the $m$ dimensions of $\mathbf{X}$ are considered independent the joint c.d.f. $F_{\mathbf{X}}(\mathbf{x})$ factorizes, and we then have by Eq. 9 that the coverage constraint in Eq. 10 reduces to
$C^{\perp}_{\mathbf{X}}(q_{*},\dots,q_{*})=F_{\mathbf{X}}(\mathbf{x}_{*})=\prod_{k=1}^{m}F_{X_{k}}(x_{*,k})=\prod_{k=1}^{m}F_{X_{k}}(F_{X_{k}}^{-1}(q_{*}))=\prod_{k=1}^{m}q_{*}=(q_{*})^{m}\geq 1-\alpha,$ (14)
where we refer to $C^{\perp}_{\mathbf{X}}$ as the factorized ‘independent copula’ (Sklar, 1973; Messoudi et al., 2021). max-rank solves the constraint by Prop. 1 and is optimally tight, thus the solution under independence $q^{\perp}_{\max}$ directly satisfies $q^{\perp}_{\max}\geq(1-\alpha)^{1/m}$. In contrast, the Bonferroni correction ensures the condition is met by $q_{\text{Bonf}}=(1-\alpha/m)$, since $(1-\alpha/m)\geq(1-\alpha)^{1/m}$ for any $m\geq 1,\,\alpha\in(0,1)$. It directly follows that $q^{\perp}_{\max}\leq q_{\text{Bonf}}$, and max-rank therefore performs at least as well as the Bonferroni correction under independence.
Under positive dependency. In this setting we assume that the components $X_{1},\dots,X_{m}$ of $\mathbf{X}$ exhibit a positive correlation known as positive lower orthant dependency, which is defined as follows:
Definition 1 (Positive lower orthant dependency (PLOD), (Nelsen, 2006), Def. 5.7.1).
An $m$-dimensional random vector $\mathbf{X}=(X_{1},\dots,X_{m})$ is PLOD if for all $\mathbf{x}=(x_{1},\dots,x_{m})\in\mathbb{R}^{m}$ we have that
$\mathbb{P}(\mathbf{X}\leq\mathbf{x})\geq\prod_{k=1}^{m}\mathbb{P}(X_{k}\leq x_{k}).$
That is, the condition supposes that the probability of components jointly taking on large (or small) values is at least as large as if all components were considered independent. PLOD provides a sensible notion of positive dependency that becomes apparent through rank ordering, and can thus be leveraged by max-rank. If we compare the copula $C_{\mathbf{X}}$ under PLOD to the independent copula $C^{\perp}_{\mathbf{X}}$ evaluated at $\mathbf{q_{*}}$, it directly follows from Def. 1 that
$C_{\mathbf{X}}(q_{*},\dots,q_{*})\geq C^{\perp}_{\mathbf{X}}(q_{*},\dots,q_{*})\geq 1-\alpha,$ (15)
i.e., $C_{\mathbf{X}}$ stochastically dominates $C^{\perp}_{\mathbf{X}}$. Thus a solution $q_{\max}$ to Eq. 10 under PLOD will be at most as large as a solution to the constraint under independence (Eq. 14), i.e., we have $q_{\max}\leq q^{\perp}_{\max}$ using max-rank. In contrast the Bonferroni correction only becomes increasingly conservative under dependency (see, e.g., Fig. 4), thus no advantage is gained over the solution under independence $q_{\text{Bonf}}$. Together with $q^{\perp}_{\max}\leq q_{\text{Bonf}}$ this implies that max-rank always performs on par or better than the Bonferroni correction, both under settings of independence and positive dependency.
We summarize our arguments by restating our proposition from § 4, which follows from the above:
Proposition 2.
The solution $q_{\max}$ provided by the max-rank procedure (Algo. 2) to Eq. 6 ensures that we have $q_{\max}\leq q^{\perp}_{\max}\leq q_{\text{Bonf}}$ in the case of both independent and positively dependent test statistics.
A.4 Using max-rank for conformal prediction
Two key conditions are imposed on any multiple testing correction that integrates with conformal prediction: (1) the exchangeability condition on the underlying data is preserved, and (2) Eq. 7 is corrected in such a way that the conformal coverage guarantee is upheld. We next show that max-rank satisfies both conditions by looking at its two key operations, namely (1) operating in the rank space of nonconformity scores, and (2) applying a $\max$-operation to these ranks. We first address the exchangeability requirement, followed by validity.
Preserving exchangeability. Our arguments rely on work by Kuchibhotla (2020), which establishes connections between exchangeability and rank-based hypothesis testing. Recall our rank definition in Eq. 5. For operating in the rank space of nonconformity scores, we make use of the following theorem:
Theorem 1 (Distribution of ranks (Kuchibhotla, 2020), Thm. 2).
For exchangeable random variables $X_{1},\dots,X_{n}$ we have that
$(\mathrm{rank}(X_{i};\{X_{1:n}\}):i\in[n])\sim\mathrm{Unif}(\{\pi:[n]\rightarrow[n]\}),$
where $\mathrm{Unif}$ is the uniform distribution over all permutations of $[n]$, i.e., each permutation occurs with equal probability $1/n!$.
It follows that the set of ranks for any underlying set of exchangeable nonconformity scores $S=\{s_{1},\dots,s_{n}\}$ is exchangeable itself, since the ranks are uniformly distributed and do not depend on the distribution of scores. We can leverage Thm. 2 below to show that particular choices of score functions ensure exchangeable score sets, but here refer to the broader fact that the score set is generally known to be exchangeable since most score functions (e.g., absolute residual scores of the form $|Y_{i}-\hat{f}(X_{i})|$) are symmetric and applied point-wise per (calibration) sample (Angelopoulos et al., 2023; Tibshirani et al., 2019; Fisch et al., 2022).
For the exchangeability preservation of the $\max$-operation, we now explicitly make use of the following theorem:
Theorem 2 (Exchangeability under transformations (Kuchibhotla, 2020; Commenges, 2003)).
Given a vector of exchangeable random variables $X=(X_{1},\dots,X_{m})\in\mathcal{X}^{m}$ and a fixed transformation $G$, we consider $G$ exchangeability-preserving if for each permutation $\pi_{1}:[m]\rightarrow[m]$ there exists a permutation $\pi_{2}:[m]\rightarrow[m]$ s.t.
$\forall x\in\mathcal{X}^{m}:\pi_{1}G(x)=G(\pi_{2}\,x).$
Following Algo. 2, we denote by $\mathbf{r}_{i}=(r_{i,1}\cdots r_{i,m})\in\mathbb{N}^{m}$ the $i$-th row of the matrix of rank statistics $\mathbf{R}$. The ranks are exchangeable by Thm. 1, and they are random variables since any data randomness propagates through the fixed score function onto the score ranks. If we fix the transformation $G$ to be the $l^{\infty}$-norm, then it follows for its application to any row $\mathbf{r}_{i},\,i\in[n]$ that
$\pi_{1}G(\mathbf{r}_{i})=\pi_{1}\lVert\mathbf{r}_{i}\rVert_{\infty}=\pi_{1}\underset{1\leq k\leq m}{\max}r_{i,k}=\underset{1\leq k\leq m}{\max}\pi_{2}r_{i,k}=\lVert\pi_{2}\mathbf{r}_{i}\rVert_{\infty}=G(\pi_{2}\,\mathbf{r}_{i}),$ (16)
satisfying Thm. 2 and the exchangeability of the $\max$-operation. Intuitively, the $l^{\infty}$-norm is symmetric and permutation-invariant, that is, a permutation $\pi_{2}$ to its arguments does not change the output permuted by $\pi_{1}$.
Ensuring validity. We next address the validity requirement, similarly tackling the two key operations of max-rank. To show that validity and thus Eq. 2 is maintained also when operating in the rank space, we employ the following corollary:
Corollary 1 ((Kuchibhotla, 2020), Corr. 1).
Under assumptions of Thm. 1 we have that
$\mathbb{P}(\text{rank}(X_{n};\{X_{1:n}\})\leq t)=\frac{\lfloor t\rfloor}{n}$
for $t\in\mathbb{N}$. In addition, the random variable $\text{rank}(X_{n};\{X_{1:n}\})/n$ is a valid p-value.
Consider a one-dimensional test sample $(Y_{n+1},X_{n+1})$, or equivalently, fix a specific $k$-th test dimension. Then we denote the extended set of nonconformity scores by $S^{\prime}=S\cup\{s_{n+1}\}$, where $s_{n+1}$ is the test sample’s score. Using Eq. 5 and a sorted score set $S^{\prime}$ observe that $\text{rank}(s_{i};S^{\prime})=i$, i.e., the rank of the $i$-th score is its position index $i$. By Eq. 4 we have that $Y_{n+1}\in\hat{C}(X_{n+1})$ is equivalent to $s_{n+1}\leq\hat{Q}(1-\alpha;S)$, which in the rank space and using the conformal quantile’s definition translates to $\text{rank}(s_{n+1};S^{\prime})\leq t$, with $t=\left\lceil(n+1)(1-\alpha)\right\rceil$. Using rank exchangeability and Cor. 1 it follows that
$\begin{split}\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1}))=\mathbb{P}(\text{rank}(s_{n+1};S^{\prime})\leq t)=\frac{\lfloor t\rfloor}{n+1}=\frac{\lceil(n+1)(1-\alpha)\rceil}{n+1}\geq\frac{(n+1)(1-\alpha)}{n+1}=1-\alpha,\end{split}$ (17)
thus ensuring target coverage $(1-\alpha)$. Alternatively, note that valid coverage also directly follows from max-rank’s FWER-controlling property (Prop. 1).
For the $\max$-operation, we make a simple argument on stochastic dominance of the quantiles. By definition of the composite empirical null $\mathbf{r}_{\max}$ in Algo. 2 we have for any $k\in[m]$ that $\hat{Q}(1-\alpha;\mathbf{r}_{\max})\geq\hat{Q}(1-\alpha;\mathbf{r}_{k})$ holds. Clearly, a prediction set $\hat{C}_{k}$ constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{k})$ is then a subset of $\hat{C}_{\max}$ constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$, i.e., $\hat{C}^{\max}\supseteq\hat{C}^{k}$. Since by Eq. 4 and validity of operating in the rank space $\hat{C}_{k}$ is constructed to ensure target coverage, the same coverage guarantee applies to $\hat{C}_{\max}$.
We summarize our results by restating our proposition from § 6.1, which follows from the above:
Proposition 3.
The max-rank procedure (Algo. 2) provides a multiple testing correction for conformal prediction that preserves exchangeability and ensures valid coverage (Eq. 2).
Appendix B ALGORITHMIC DETAILS
Algorithm 3 Multiple testing correction via max-T (Westfall and Young (1993))
1:Input: Hypothesis tests $k=1,\dots,m$ and associated test statistics $(T_{1},\dots,T_{m})\in\mathbb{R}^{m}$, a resampling counter $i\in[n]$, and empty vector of resampled test statistics $\mathbf{t}_{\max}\in\mathbb{R}^{n}$.
2:Output: Adjusted p-values $(P^{\text{adj}}_{1},\dots,P^{\text{adj}}_{m})\in\mathbb{R}^{m}$ with FWER control at level $\alpha$.
3:Procedure:
4:For $i=1,\dots,n$: Resample data, e.g., via permutation or bootstrapping. Recompute test statistics for $i$-th resampling and store maximum statistic: $\mathbf{t}_{\max,i}=\max_{1\leq k\leq m}T_{i,k}$.
5:Compute adjusted p-value by comparing observed statistics to the maximum statistics under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\mathbf{t}_{\max,i}\geq T_{k}]$.
6:Decision rule: $\forall k\in[m]$, we cannot reject the null if $P^{\text{adj}}_{k}>\alpha$.
7:End procedure
Algorithm 4 Multiple testing correction via min-P (Westfall and Young (1993))
1:Input: Hypothesis tests $k=1,\dots,m$ and associated test statistics $(T_{1},\dots,T_{m})\in\mathbb{R}^{m}$, a resampling counter $i\in[n]$, and empty vector of resampled p-values $\mathbf{p}_{\min}\in[0,1]^{n}$.
2:Output: Adjusted p-values $(P^{\text{adj}}_{1},\dots,P^{\text{adj}}_{m})\in[0,1]^{m}$ with FWER control at level $\alpha$.
3:Procedure:
4:For all test statistics, compute the associated unadjusted p-values $(P_{1},\dots,P_{m})\in[0,1]^{m}$.
5:For $i=1,\dots,n$: Resample data, e.g., via permutation or bootstrapping. Recompute p-values for $i$-th resampling and store the minimum p-value: $\mathbf{p}_{\min,i}=\min_{1\leq k\leq m}P_{i,k}$.
6:Compute adjusted p-value by comparing observed p-values to the minimum p-values under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\mathbf{p}_{\min,i}\leq P_{k}]$.
7:Decision rule: $\forall k\in[m]$, we cannot reject the null if $P^{\text{adj}}_{k}>\alpha$.
8:End procedure
The algorithmic descriptions for Westfall and Young (1993)’s max-T and min-P multiple testing corrections are given in Algo. 3 and Algo. 4 respectively. The main paper contains the algorithms for split conformal prediction in the general case (Algo. 1) and our correction max-rank (Algo. 2).
Appendix C ADDITIONAL BACKGROUND
C.1 FDR control
In contrast to the stronger notion of FWER control which bounds the probability of committing at least one Type-I error, controlling the false discovery rate (FDR) targets a limit on the number of false discoveries in expectation across hypothesis tests. This weaker notion of Type-I error control comes at the benefit of a larger fraction of significant discoveries (Shaffer, 1995). Denote the number of false positives $m_{\text{FP}}$ as before, and the fraction of correct discoveries (i.e., true positives) as $m_{\text{TP}}$. Then, FDR control at significance level $\alpha$ ensures that
$\mathbb{E}_{m}(m_{\text{FP}}/(m_{\text{FP}}+m_{\text{TP}}))\leq\alpha,$ (18)
where the expectation is taken over the $m$ hypotheses. FDR control is appropriate in settings in which a limited proportion of false positives can be tolerated in exchange for a greater chance at significant results, such as in genomics studies (Han et al., 2009). Commonly known methods for FDR control include the Benjamini-Hochberg (Benjamini and Hochberg, 1995), Benjamini-Yekutieli (Benjamini and Yekutieli, 2001) and Simes corrections (Simes, 1986) [^6] Simes was originally designed for FWER control, but then adapted to FDR control.. In contrast to classical corrections with FWER control which assume independence, FDR-controlling procedures are generally more robust to dependencies between tests.
C.2 FWER control via the Bonferroni correction
We demonstrate how the Bonferroni correction ensures FWER control using a simple union bound argument. Denote for our $m$ hypotheses the corresponding p-values as $P_{1},\dots,P_{m}$. Observe that in the main paper any permutation-based p-values (i.e., on the basis of Eq. 3) are described as $\hat{P}_{1},\dots,\hat{P}_{m}$ to highlight their data-driven nature, whereas we do not specify here how the Bonferroni p-values are computed. Let us label the fraction of hypotheses that should be correctly marked as non-significant (i.e., true negatives) as $m_{\text{TN}}$. The null is rejected when the respective p-value reaches below significance level, i.e., when we have $P_{k}\leq\alpha_{\text{Bonf}}$ for some test $k$. Evaluating the risk of incorrectly rejecting a test that belongs to $m_{\text{TN}}$ (i.e., generating a false positive), we obtain that
$\text{FWER}=\mathbb{P}\left(\bigcup_{k=1}^{m_{\text{TN}}}(P_{k}\leq\alpha_{\text{Bonf}})\right)=\mathbb{P}\left(\bigcup_{k=1}^{m_{\text{TN}}}(P_{k}\leq\frac{\alpha}{m})\right)\leq\sum_{k=1}^{m_{\text{TN}}}\mathbb{P}(P_{k}\leq\frac{\alpha}{m})=\sum_{k=1}^{m_{\text{TN}}}\frac{\alpha}{m}=\frac{m_{\text{TN}}}{m}\alpha\leq\alpha.$ (19)
We observe that since it follows that FWER $\leq\alpha$, employing $\alpha_{\text{Bonf}}=\alpha/m$ provides valid FWER control.
Appendix D ADDITIONAL EXPERIMENT DETAILS
D.1 Simulation study: FWER control and sample size
Figure 4: We examine global adjusted significance levels $\hat{\alpha}$ against the desired FWER control level $\alpha=0.05$ for varying correlations $\rho$, number of tests $m$ and sample size $n$. Empirical FWER control rates are directly related to the magnitudes of $m$ and $n$, suggesting that larger number of parallel tests require larger sample sizes. We also see that Bonferroni is consistently more conservative than max-rank across all combinations.
[…]
The observed interactions between quantities seem consistent and suggest a meaningful relationship that warrants future theoretical investigation, as mentioned in § 7. It would be desirable to quantify the expected FWER control achievable for a given significance level $\alpha$, sample size $n$ and number of tests $m$ in analytical form.
D.2 Multiple testing correction baselines
We consider the following multiple testing correction baselines in our experiments (§ 6.2): Bonferroni, Bonferroni-Šidák, Gumbel Copula, Empirical Copula, and max-T. We refer to § C.2 for the Bonferroni correction and § 3.2 as well as Algo. 3 for the max-T correction. We comment on the remaining corrections.
[…]
Gumbel Copula. The Gumbel copula is a parametric copula which, for some multivariate input $\mathbf{u}=(u_{1},\dots,u_{m})\in[0,1]^{m}$, evaluates to
$C_{\text{Gumbel}}(\mathbf{u})=\exp\left\{-\sum_{k=1}^{m}(-\ln u_{k})^{\theta}\right\}^{1/\theta}.$ (20)
It has a light-weight parametric form that only depends on a single parameter $\theta$, which is estimated numerically. Messoudi et al. (2021) show that under the restriction $u=u_{1}=\dots=u_{m}$ (i.e., $\alpha=\alpha_{1}=\dots=\alpha_{m}$ and thus $\mathbf{u}=(1-\alpha,\dots,1-\alpha)$ in our setting) a solution to the coverage condition in Eq. 6 is given analytically as $\alpha_{\text{Gumbel}}=1-(1-\alpha)^{1/\sqrt[\theta]{m}}$. Our adjusted significance level is then obtained by plugging in the estimate $\hat{\theta}$ into the equation. We implement both the Gumbel and empirical copulas using the library copulae [^7] See https://github.com/danielbok/copulae.
Empirical Copula. The empirical copula is the copula-equivalent of an empirical, non-parametric estimator of the joint cumulative distribution. It evaluates to
$C_{\text{Emp}}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\prod_{k=1}^{m}\mathbbm{1}[u_{i,k}\leq u_{k}],$ (21)
where $\mathbbm{1}[\cdot]$ is the indicator function, $u_{k}$ is a component of $\mathbf{u}$ as above (equivalent to $1-\alpha$), and $u_{i,k}$ is the cumulative probability of the $i$-th sample under its marginal c.d.f. in the $k$-th dimension (estimated empirically), i.e., $u_{i,k}=\hat{F}_{X_{k}}(x_{i,k})$. We can think of the copula as empirically estimating the fraction of samples jointly satisfying the condition $\mathbbm{1}[u_{i,k}\leq u_{k}]$ in all dimensions, since only in such cases does the product evaluate to one. As there is no analytical solution (Messoudi et al., 2021) we have to search through a grid of possible adjusted significance levels $\hat{\alpha}\in(0,1)$, evaluate them under the copula, and select the largest one satisfying target coverage $C_{\text{Emp}}(1-\hat{\alpha},\dots,1-\hat{\alpha})\geq(1-\alpha)$, thus directly addressing the problem in Eq. 6 (note that $\min q_{*}\Leftrightarrow\max\hat{\alpha}$).
D.3 Conformal prediction metrics: empirical coverage and prediction set size
Our approaches are evaluated across key metrics employed in conformal prediction which jointly assess its main desiderata (Vovk et al., 2005; Angelopoulos et al., 2023): empirical coverage to assess validity, i.e., the coverage target in Eq. 2, and prediction set size to assess efficiency. The efficiency of the solution captures its usefulness in practice, since a trivial solution such as an infinite set would meet coverage goals but be of little use. Another popular assessment criterion is adaptivity, i.e., the ability to cover equally across different subgroups of the data. Since we are less preoccupied with the nature of our conformal guarantee in this work (marginal vs. conditional vs. balanced), we do not consider it here.
Moving beyond a single test sample $(X_{n+1},Y_{n+1})$ as frequently denoted in the main paper, let us specify a test set of arbitrary size $n_{t}$ as $\mathcal{D}_{test}=\{(X_{j},Y_{j})\}_{j=n+1}^{n+n_{t}}$. We only consider a test sample covered when all its $k\in[m]$ test dimensions are jointly covered. Thus the average empirical coverage is given by
$\text{Coverage}=\frac{1}{n_{t}\,m}\sum_{j=n+1}^{n+n_{t}}\mathbbm{1}\left[\bigcap_{k=1}^{m}(Y_{j,k}\in\hat{C}(X_{j,k}))\right],$ (22)
where $\mathbbm{1}[\cdot]$ is the indicator function. We note that the empirical coverage is a random quantity parametrized by a coverage distribution, and can slightly deviate from target coverage $(1-\alpha)$ based on factors such as the size of the calibration data $|\mathcal{D}_{cal}|$ (Vovk et al., 2005). For efficiency, we compute the average prediction set size across samples and test dimensions as
Interval $\text{Interval Width}=\frac{1}{n_{t}\,m}\sum_{j=n+1}^{n+n_{t}}\sum_{k=1}^{m}|\hat{C}(X_{j,k})|.$ (23)
Note that since our regression experiments in § 6.2 produce prediction sets in the form of intervals, we directly define the metric as prediction interval width rather than the more generic term on prediction set size.
D.4 Conformal multi-target regression
We provide a few more details on our experimental design, which follows closely Messoudi et al. (2021). We select a subset of three interesting datasets with varying numbers of regression targets and sample size lengths [^8] Obtainable at https://github.com/tsoumakas/mulan: scpf (website traffic data, $m=3$, 1137 samples), rf1 (river flow data, 9125 samples, $m=8$), and scm1d (product pricing data, 9803 samples, $m=16$). We use the data pre-processing scheme of Messoudi et al. (2021) [^9] See https://github.com/M-Soundouss/CopulaConformalMTR and similarly train a multi-target random forest regressor using sci-kit learn (Pedregosa et al., 2011) with minimal hyperparameter tuning. Each dataset is randomly split into 60% training, 20% calibration, and 20% test data, and the model training and conformal procedure is repeated and results averaged across 100 random trials for robust evaluation.
D.5 Conformal object detection
Object detection tasks consist of a dual problem wherein both an object’s location, as parametrized via it’s four bounding box coordinates (thus $m=4$), and the object’s class label need to be predicted. We focus on applying conformal prediction to the bounding box prediction task, which forms a regression problem in the (approximately) continuous pixel space of input images. Our experimental design strongly relates to Mukama et al. (2024) but builds upon our own developed conformal methodology for object detection (Timans et al., 2024). Instead of considering only objects of a single class ‘car’, we instead apply conformal prediction separately to a variety of classes. This is easily achieved by filtering the labelled calibration data by class label, and computing class-specific conformal quantiles via Algo. 1 used to build prediction sets for each test sample based on its (unconformalized) class label. Thus, our results in Fig. 3 averaged across multiple classes (and trials) provide a more robust performance evaluation of each multiple testing correction.
[…]
Correction Object class Interval Width Coverage
Bonferroni Class avg. 42.9513 0.9134
Person 27.0747 0.9131
Bicycle 55.6782 0.9160
Car 37.9270 0.9135
Motorcycle 60.1595 0.9128
Bus 60.4550 0.9130
Truck 67.4933 0.9122
Traffic Light 15.8933 0.9083
Stop Sign 18.9295 0.9187
Bonf.-Šidák Class avg. 42.1049 0.9102
Person 26.5748 0.9098
Bicycle 54.7535 0.9132
Car 37.0510 0.9103
Motorcycle 59.4491 0.9103
Bus 58.9098 0.9094
Truck 65.9317 0.9089
Traffic Light 15.6576 0.9048
Stop Sign 18.5114 0.9153
Gumbel Copula Class avg. 40.444 0.9039
Person 25.8873 0.9048
Bicycle 53.1460 0.9085
Car 35.4183 0.9038
Motorcycle 57.3168 0.9015
Bus 56.4559 0.9032
Truck 62.3353 0.9014
Traffic Light 15.3669 0.9006
Stop Sign 17.6255 0.9070
Emp. Copula Class avg. 39.9836 0.9015
Person 25.2839 0.8998
Bicycle 51.5578 0.9028
Car 34.5660 0.9001
Motorcycle 58.2203 0.9053
Bus 55.6454 0.9009
Truck 62.1119 0.9009
Traffic Light 15.3653 0.9006
Stop Sign 17.1182 0.9013
max-T Class avg. 41.1136 0.9002
Person 25.2977 0.8996
Bicycle 54.9369 0.9007
Car 34.6362 0.8999
Motorcycle 61.7477 0.8988
Bus 57.5168 0.9005
Truck 62.0357 0.9007
Traffic Light 15.3196 0.9005
Stop Sign 17.4184 0.9008
max-rank (Ours) Class avg. 39.6185 0.8999
Person 25.2571 0.8996
Bicycle 51.0433 0.9005
Car 34.5276 0.8999
Motorcycle 56.5972 0.8988
Bus 55.3103 0.9000
Truck 61.8884 0.9005
Traffic Light 15.3437 0.9003
Stop Sign 16.9807 0.8994
Table 2: Empirical coverage and mean prediction interval width for different corrections ($\alpha=0.1$) on BDD100k for individual classes and the class average. Results are averaged across 100 random trials. max-rank consistently provides strong results also on a per-class basis (best results, second best). We have varying object counts per class, equivalent to the number of per-class calibration samples $n$. By class from top to bottom: 44’960 (avg.), 35’748, 2032, 280’277, 883, 5151, 10’975, and 21’950.
Figure 5: A visual example of the constructed conformal prediction intervals with max-rank on a test image for the class ‘car’ (and not ‘truck’). True bounding boxes are in blue, two-sided prediction interval regions are shaded in orange.

### v4

@@
 Abstract
-Multiple hypothesis testing (MHT) commonly arises in various scientific fields, from genomics to psychology, where testing many hypotheses simultaneously increases the risk of Type-I errors.
-These errors can mislead scientific inquiry, rendering MHT corrections essential.
-In this paper, we address MHT within the context of conformal prediction, a flexible method for predictive uncertainty quantification.
-Some conformal prediction settings can require simultaneous testing, and positive dependencies among tests typically exist.
-We propose a novel correction named max-rank that leverages these dependencies, whilst ensuring that the joint Type-I error rate is efficiently controlled.
-Inspired by permutation-based corrections from Westfall and Young (1993), max-rank exploits rank order information to improve performance, and readily integrates with any conformal procedure.
-We demonstrate both its theoretical and empirical advantages over the common Bonferroni correction and its compatibility with conformal prediction, highlighting the potential to enhance predictive uncertainty quantification.
+Multiple hypothesis testing (MHT) frequently arises in scientific inquiries, and concurrent testing of multiple hypotheses inflates the risk of Type-I errors or false positives, rendering MHT corrections essential.
+This paper addresses MHT in the context of conformal prediction, a flexible framework for predictive uncertainty quantification.
+Some conformal applications give rise to simultaneous testing, and positive dependencies among tests typically exist.
+We introduce max-rank, a novel correction that exploits these dependencies whilst efficiently controlling the family-wise error rate.
+Inspired by existing permutation-based corrections, max-rank leverages rank order information to improve performance and integrates readily with any conformal procedure.
+We establish its theoretical and empirical advantages over the common Bonferroni correction and its compatibility with conformal prediction, highlighting the potential to strengthen predictive uncertainty estimates.
 1 INTRODUCTION
-Multiple hypothesis testing (MHT) occurs whenever we attempt to draw conclusions over multiple statistical hypothesis at once (Shaffer, 1995).
-Such scenarios frequently occur in various scientific domains, ranging from medicine and genomics (Ludbrook, 1994; Qu et al., 2010; Storey and Tibshirani, 2003) to economics and psychology (Blakesley et al., 2009; List et al., 2019).
-A crucial issue with MHT is that as the number of tested hypotheses increases, so does the likelihood of committing a Type-I error, i.e., erroneously rejecting a null in favour of significance.
-Such false positives may lead scientists astray by drawing incorrect conclusions and pursuing costly but statistically meaningless studies (Benjamini, 2010).
-Thus, a large body of research in statistics has preoccupied itself with multiple testing corrections to permit valid statistical inference even in such scenarios (Bender and Lange, 2001).
-In this work, we investigate a particular instance of the multiple testing problem: predictive uncertainty quantification via conformal prediction (Shafer and Vovk, 2008; Vovk et al., 2005).
-Conformal prediction is an efficient way to produce statistically rigorous uncertainties for complex predictive models, such as deep neural networks (Angelopoulos et al., 2023).
-MHT issues can arise from the need to run multiple conformal procedures at once (e.g., in multi-target regression (Messoudi et al., 2021)), necessitating a correction.
-Moreover, these conformal procedures are known to exhibit positive dependency (Bates et al., 2022), rendering simple corrections such as Bonferroni inefficient a priori.
-Exploiting the natural connections between conformal prediction and permutation testing, we propose a simple MHT correction that is able to exploit the positive dependency between hypotheses when apparent through rank ordering.
-Furthermore, it ensures strong Type-I error control on the family-wise error rate (FWER), i.e., the probability of a single false discovery.
-Our approach is closely connected to the permutation-based corrections of Westfall and Young (1993), and readily integrates with the conditions assumed by conformal prediction.
-To summarize, our contributions include:
-  • Proposing an efficient correction called max-rank for MHT settings with positive dependency (Algo.
-2) and establishing its connection to the correction procedures of Westfall and Young (1993) (§ 3.2).
-  • Proving that max-rank is superior to the Bonferroni correction in terms of FWER control via multivariate copula arguments (§ 4, § 5).
-  • Demonstrating both theoretically (§ 6.1) and empirically (§ 6.2) max-rank’s suitability for conformal prediction.
+Multiple hypothesis testing (MHT) occurs whenever we attempt to draw conclusions over multiple statistical hypotheses at once (Shaffer, 1995).
+Such scenarios regularly occur across different scientific domains ranging from medicine and genomics (Ludbrook, 1994; Qu et al., 2010; Storey and Tibshirani, 2003) to economics and psychology (Blakesley et al., 2009; List et al., 2019).
+A key issue is that as the number of tests grows, so does the risk of Type-I errors – falsely rejecting the null hypothesis – which can mislead scientific inquiry and waste resources on spurious or incorrect findings (Benjamini, 2010).
+To mitigate this, a vast body of statistical research has focused on multiple testing corrections that permit valid inference even in such cases (Bender and Lange, 2001).
+In this work, we examine a specific instance of the multiple testing problem: predictive uncertainty quantification through conformal prediction (Shafer and Vovk, 2008; Vovk et al., 2005).
+Conformal prediction offers a statistically rigorous framework for uncertainty estimation in complex predictive models, such as deep neural networks (Angelopoulos et al., 2023).
+However, MHT challenges arise when multiple conformal procedures are run simultaneously (e.g., for multi-target regression (Messoudi et al., 2021)), necessitating a correction.
+Moreover, these procedures often exhibit positive dependencies (Bates et al., 2022), rendering standard approaches like Bonferroni inefficient a priori.
+footnotetext: Corresponding author: $<$a.r.timans@uva.nl$>$footnotetext: ${}^{\text{\textdagger}}$ignificant contributions.
+Leveraging the natural connections between conformal prediction and permutation testing, we introduce a simple MHT correction that exploits positive dependencies among hypotheses when evident through rank ordering.
+Our method ensures strong control of the family-wise error rate (FWER), i.e., the probability of at least one false discovery.
+Closely related to the permutation-based corrections of Westfall and Young (1993), it readily integrates with the conditions assumed by conformal prediction without imposing any additional assumptions.
+In summary, we contribute:
+  • max-rank, an efficient correction for MHT settings with positive dependency (Algo. 2), and establish its connection to the multiple testing procedures of Westfall and Young (1993) (§ 3.2).
+  • Prove that max-rank offers superior FWER control over the Bonferroni correction using multivariate copula arguments (§ 4, § 5).
+  • Demonstrate both theoretically (§ 6.1) and empirically (§ 6.2) max-rank’s suitability for conformal prediction settings with multiple testing.
 2 BACKGROUND & RELATED WORK
-We begin by commenting on our multiple testing setting and forms of Type-I error control, with a closer look at the Bonferroni correction.
-We then thoroughly introduce the conformal prediction setting and its connection to hypothesis testing.
+We first discuss our multiple testing setting and Type-I error control, with a closer look at the Bonferroni correction.
+Then, we introduce the conformal prediction setting and its connection to hypothesis testing.
 Multiple testing in parallel and in series.
@@
 In contrast, testing in series suggests conclusions are drawn one after another, and the outcome of a test may determine whether and which subsequent hypotheses should be explored.
-For example, a clinical drug trial with multiple phases may only initiate a follow-up if the results are satisfactory during the first phase (Bauer, 1991).
+For example, a clinical drug trial with multiple phases may only initiate a follow-up if results are satisfactory during the first phase (Bauer, 1991).
 Family-wise error rate control.
-We may distinguish two forms of controlling the Type-I error rate during testing: control in the strong sense via the family-wise error rate (FWER), and in the weaker sense via the false discovery rate (FDR, see § C.1).
-FWER control, the more stringent of the two, aims to limit the probability of committing at least one Type-I error across all considered hypotheses.
+We can further distinguish among different types of Type-I error control while testing: control in the strong sense via the FWER, and in weaker forms via, e.g., the false discovery rate (see § C.1).
+FWER control as the most stringent form aims to limit the probability of committing at least one Type-I error across all considered hypotheses.
 If we test $m$ hypotheses and denote by $m_{\text{FP}}$ the number of falsely rejected null hypotheses (i.e., false positives), then FWER control at significance level $\alpha$ ensures that
 $\mathbb{P}(m_{\text{FP}}\geq 1)\leq\alpha.$ (1)
-Such strict error control is appropriate in settings where false discoveries may have strong implications, such as in drug testing and clinical trials.
-Since conformal prediction is oftentimes motivated from a safety perspective by its provided coverage guarantees (see § 2.1) and can be applied to similar settings (e.g., Lei and Candès (2021); Candès et al.
+This strict error control is appropriate in settings where false discoveries may have strong implications, such as in drug testing or clinical trials.
+Since conformal prediction is oftentimes motivated from a safety perspective by its provided coverage guarantees and is applicable to such settings (e.g., Lei and Candès (2021); Candès et al.
 (2023); Yin et al.
-(2024)), FWER control is also an appropriate notion here.
+(2024)), FWER control is also the appropriate notion here.
 FWER control via the Bonferroni correction.
@@
 Albeit MHT issues have been addressed as early as Tukey (1953), the research on valid and efficient corrections under dependency remains remarkably active.
-Westfall and Young (1993) suggested an early resampling-based procedure for capturing data dependencies, while recent years have seen further work on arbitrary dependencies (Causeur et al., 2009; Sun and Tony Cai, 2009; Leek and Storey, 2008; Liu and Xie, 2019; Finner et al., 2017; Gou and Tamhane, 2018; Chi et al., 2022) and p-value merging (Wilson, 2019; Vovk et al., 2022; Chen et al., 2023).
+Westfall and Young (1993) suggested an early resampling-based procedure for capturing data dependencies, while recent years have seen further work on arbitrary dependencies (Causeur et al., 2009; Sun and Tony Cai, 2009; Leek and Storey, 2008; Liu and Xie, 2019; Finner et al., 2017; Gou and Tamhane, 2018; Chi et al., 2025) and p-value merging (Wilson, 2019; Vovk et al., 2022; Chen et al., 2023).
 Indeed, a remark by Chi et al.
-(2022) that “the gap between theory and practice is mildly unsettling” suggests these problems remain far from fully solved.
+(2025) that “the gap between theory and practice is mildly unsettling” suggests these questions remain far from fully solved.
 Multiple testing in conformal prediction.
-Conformal applications necessitating some form of multiple testing correction include multi-step time series (Stankeviciute et al., 2021; Sun and Yu, 2023; Xu et al., 2024) and trajectory estimation (Lindemann et al., 2023; Muthali et al., 2023); multi-target classification (Cauchois et al., 2021; Fisch et al., 2021) and regression (Messoudi et al., 2021, 2022; Feldman et al., 2023), including conformal object detection (Andéol et al.
+Conformal applications necessitating some form of multiple testing correction are widespread, and include multi-step time series and trajectory estimation (Stankeviciute et al., 2021; Sun and Yu, 2023; Xu et al., 2024; Lindemann et al., 2023; Muthali et al., 2023); multi-target classification and regression (Cauchois et al., 2021; Fisch et al., 2021; Messoudi et al., 2021, 2022; Feldman et al., 2023), among them conformal object detection (Andéol et al.
 (2023); Mukama et al.
-(2024), see § 6.2); treatment analysis (Lei and Candès, 2021; Yin et al., 2024) and screening (Jin and Candès, 2023; Fisch et al., 2022), or even accelerated inference (Schuster et al., 2021).
-Throughout, employed corrections range from Bonferroni (Stankeviciute et al., 2021; Lindemann et al., 2023; Muthali et al., 2023; Andéol et al., 2023; Jin and Candès, 2023) to Copulas (Messoudi et al., 2021; Sun and Yu, 2023), Simes (Fisch et al., 2021) and other forms of sequential testing (Lei and Candès, 2021; Yin et al., 2024; Cauchois et al., 2021; Angelopoulos et al., 2021), showcasing that a variety of solutions have been considered.
-2.1 Conformal prediction
+(2024), see § 6.2); treatment analysis and screening (Lei and Candès, 2021; Yin et al., 2024; Jin and Candès, 2023; Fisch et al., 2022), and even as diverse as early-exiting and diffusion models (Schuster et al., 2021; Teneggi et al., 2023).
+Throughout, employed MHT corrections frequently include Bonferroni (Stankeviciute et al., 2021; Lindemann et al., 2023; Muthali et al., 2023; Andéol et al., 2023; Jin and Candès, 2023), but can range from Copulas (Messoudi et al., 2021; Sun and Yu, 2023) to Simes (Fisch et al., 2021) and other forms of sequential testing (Lei and Candès, 2021; Yin et al., 2024; Cauchois et al., 2021; Angelopoulos et al., 2021), showcasing a variety of considered solutions.
+Naturally, the application-specific test design (in parallel vs.
+in series) largely governs the suitability of employing particular corrections.
 Algorithm 1 Split conformal prediction (general case)
-1:Input: Pre-trained predictor $\hat{f}$, calibration data $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$, test sample $(X_{n+1},Y_{n+1})\sim\mathcal{P}_{XY}$, miscoverage rate $\alpha\in(0,1)$.
+1:Input: Pre-trained predictor $\hat{f}$, calibration samples $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$, test sample $(X_{n+1},Y_{n+1})\sim\mathcal{P}_{XY}$, miscoverage rate $\alpha\in(0,1)$.
 2:Output: Prediction set $\hat{C}(X_{n+1})$ for test sample.
@@
 […]
-6:For the new test sample $(X_{n+1},Y_{n+1})$, a valid prediction set for $X_{n+1}$ is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:\,s(\hat{f}(X_{n+1}),y)\leq\hat{Q}(1-\alpha;S)\}$.
-Validity refers to satisfying Eq.
-2.
+6:For a new test sample $(X_{n+1},Y_{n+1})$, a valid prediction set satisfying Eq. 2 is given by $\hat{C}(X_{n+1})=\{y\in\mathcal{Y}:\,s(\hat{f}(X_{n+1}),\,y)\leq\hat{Q}(1-\alpha;S)\}$.
 7:End Procedure
-Conformal prediction (CP) is a distribution-free, post-hoc, and finite-sample approach to predictive uncertainty quantification (Vovk et al., 2005; Shafer and Vovk, 2008).
-The obtained uncertainty around a predictor’s outputs is expressed in terms of prediction sets, and provides an associated guarantee on coverage (Eq.
+Conformal prediction.
+Conformal prediction (CP) offers a distribution-free, post-hoc, and finite-sample approach to predictive uncertainty quantification (Vovk et al., 2005; Shafer and Vovk, 2008).
+The obtained uncertainty around a predictor’s output is expressed as a prediction set, and provides an associated guarantee on coverage (Eq.
 2).
@@
 1.
-Split CP requires a separate calibration dataset $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$ on which dissimilarity measures (so-called nonconformity scores) between predicted and true values are computed.
-For a new test sample $(X_{n+1},Y_{n+1})$, the prediction set is then constructed by including any predictions that align with the previously observed scores on $\mathcal{D}_{cal}$, as determined by a conformal quantile $\hat{Q}_{1-\alpha}$ that ensures a coverage rate of $(1-\alpha)$ for true values.
+Split CP requires a separate calibration dataset $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim\mathcal{P}_{XY}$ for which dissimilarity values (so-called nonconformity scores) between predicted and true outputs are computed.
+Given a new test sample $(X_{n+1},Y_{n+1})\sim\mathcal{P}_{XY}$, a prediction set is then constructed by including any predictions that align with the previously observed scores on $\mathcal{D}_{cal}$, as governed by a conformal quantile $\hat{Q}_{1-\alpha}$ ensuring a coverage rate of $(1-\alpha)$ for true values.
 The approach is distribution-free since it does not impose any distributional assumptions.
-Instead, the key requirement is the exchangeability of $\mathcal{D}_{cal}\,\cup\,\{(X_{n+1},Y_{n+1})\}$, a relaxed i.i.d.
-assumption suggesting permutation invariance of the samples.
-It is post-hoc because it can be used with any ‘black-box’ pre-trained predictor without the need for training access.
-And it is finite-sample because the obtainable coverage guarantees do not rely on asymptotics.
-Specifically, if the conditions of CP are met, a probabilistic statement on containment of $Y_{n+1}$ in the prediction set $\hat{C}(X_{n+1})$ for some tolerated miscoverage rate $\alpha\in(0,1)$ is given as
+Instead, a key requirement is the exchangeability of $\mathcal{D}_{cal}\,\cup\,\{(X_{n+1},Y_{n+1})\}$, a relaxed i.i.d.
+assumption suggesting permutation invariance of samples.
+It is post-hoc because it can be used with any ‘black-box’ pre-trained predictor without the need for training or internal access.
+And it is finite-sample because the obtainable coverage guarantee does not rely on asymptotics.
+Specifically, if CP’s conditions are met, a probabilistic statement on inclusion of $Y_{n+1}$ in the prediction set $\hat{C}(X_{n+1})$ for a tolerated miscoverage rate $\alpha\in(0,1)$ is given as
 $\mathbb{P}(Y_{n+1}\in\hat{C}(X_{n+1}))\geq 1-\alpha.$ (2)
-Importantly, the provided coverage guarantee is marginally valid, i.e., it only holds on average across test and calibration data and may thus fail to cover for any particular observation.
+Importantly, the provided coverage guarantee is (i) marginally valid, i.e., it only holds on average across test and calibration data and may thus fail to cover for any particular observation; and (ii) fails to appropriately cover if exchangeability is violated, such as under strong distribution shifts (Barber et al., 2023).
 Conformal prediction as hypothesis testing.
@@
 1 can also be interpreted from a hypothesis testing perspective (Shi et al., 2013; Bates et al., 2022; Vovk et al., 2005).
-Given a prediction set candidate value $y\in\mathcal{Y}$, a statistical hypothesis for the test sample is established as $H_{0}:Y_{n+1}=y,\,\,H_{1}:Y_{n+1}\neq y$, with test statistic $s_{n+1}=s(\hat{f}(X_{n+1}),y)$ equal to its nonconformity score.
-The set of calibration scores $S=\{s_{1},\dots,s_{n}\}$ provides an empirical null distribution to compute a p-value $\hat{P}_{n+1}(y;S)$ as
-$\hat{P}_{n+1}(y;S)=|\{i=1,\dots,n:s_{i}\geq s_{n+1}\}|/(n+1),$ (3)
+Given a candidate value $y\in\mathcal{Y}$, a statistical test for its inclusion in the prediction set can be formalized by the hypothesis pair $H_{0}:Y_{n+1}=y,\,\,H_{1}:Y_{n+1}\neq y$, with test statistic $s_{n+1}=s(\hat{f}(X_{n+1}),y)$ equating its nonconformity score.
+The set of calibration scores $S=\{s_{1},\dots,s_{n}\}$ provides an empirical null distribution to compute a rank-based p-value $\hat{P}_{n+1}(y;S)$ as
+$\hat{P}_{n+1}(y;S)=\frac{|\{i=1,\dots,n+1:s_{i}\geq s_{n+1}\}|}{n+1},$ (3)
 i.e., the fraction of samples which conform worse than $s_{n+1}$.
 A testing decision on not rejecting $H_{0}$ – that is, inclusion in the prediction set – is taken at significance level $\alpha$ as $\hat{P}_{n+1}(y;S)>\alpha$.
-This approach can be interpreted as permutation testing: we are assessing if the observed nonconformity for our candidate value does not stray too much from previous observations, and the sample $(X_{n+1},y)$ thus permutes with $D_{cal}$.
-We may then interpret running multiple conformal procedures at once as permutation testing in parallel, typically at identical significance level $\alpha$ across tests.
+The formulated testing approach translates to an instance of permutation testing: we are assessing if the observed nonconformity for candidate value $y$ does not stray overly far from previous observations, and the sample $(X_{n+1},y)$ thus permutes with $D_{cal}$.
+Subsequently, we interpret running multiple conformal procedures at once as permutation testing in parallel, and at identical level $\alpha$ across tests.
 Equivalence of events.
 By the prediction set construction outlined in Algo.
-1 and the above testing perspective, a useful equivalence of events for some set candidate value $y\in\mathcal{Y}$ can be established as
+1 and the above testing perspective, a useful equivalence of events for a prediction set candidate value $y$ can be established as
 $y\in\hat{C}(X_{n+1})$ $\Leftrightarrow s_{n+1}\leq\hat{Q}(1-\alpha;S)\Leftrightarrow\hat{P}_{n+1}(y;S)>\alpha,\text{}$
 $y\notin\hat{C}(X_{n+1})$ $\Leftrightarrow s_{n+1}>\hat{Q}(1-\alpha;S)\Leftrightarrow\hat{P}_{n+1}(y;S)\leq\alpha.$ (4)
-Notably, the conformal quantile $\hat{Q}_{1-\alpha}$ [^1] We use $\hat{Q}_{1-\alpha}$ as a shortened notation for $\hat{Q}(1-\alpha;S)$.
-ensuring a fraction of $(1-\alpha)$ values below it is the equivalent of a permutation test at significance level $\alpha$.
-In both cases, the Type-I error is controlled at the same level $\alpha$.
+Notably, determining set inclusion via the conformal quantile $\hat{Q}(1-\alpha;S)$ (abbreviated to $\hat{Q}_{1-\alpha}$) which ensures a coverage rate of $(1-\alpha)$ matches a permutation test using $\hat{P}_{n+1}(y;S)$ at significance level $\alpha$.
+In both cases, Type-I error control at the same rate $\alpha$ is achieved.
 We further comment on the quantile’s properties in the following remarks:
@@
 The smallest possible quantile $\hat{Q}_{1-\alpha}$ which maintains Type-I error control at level $\alpha$ is tight in terms of coverage, i.e., it provides the smallest and thus most efficient conformal prediction sets.
+We may loosely interpret this as the matching permutation test ensuring highest statistical power.
 3 MULTIPLE TESTING CORRECTION VIA MAX-RANK
@@
 Notation.
-For consistency, we denote matrices as upper-case bolded ($\mathbf{M}$), vectors as lower-case bolded ($\mathbf{v}$), scalars without bolding ($s$).
-A multivariate random variable may also be upper-case bolded given the context ($\mathbf{X}$).
+Matrices are denoted as upper-case bold ($\mathbf{M}$), vectors as lower-case bold ($\mathbf{v}$), scalars without bolding ($s$).
+A multivariate random variable may also be upper-case bold given the context ($\mathbf{X}$).
 An index set $\{1,\dots,n\}$ is abbreviated $[n]$ for any $n\geq 1$.
 Similarly, a sequence $x_{1},\dots,x_{n}$ is shortened to $x_{1:n}$.
-The rank of an element $x_{i}$ in a set $\{x_{1:n}\}$ of $n$ distinct elements is given by
+The rank of element $x_{i}$ in a set $\{x_{1:n}\}$ of $n$ distinct elements (potentially by introducing jitter) is given by
 […]
@@
 Finally, the $l^{\infty}$-norm is given by $\lVert\mathbf{x}\rVert_{\infty}=\max_{1\leq i\leq n}|x_{i}|$ for a discrete vector $\mathbf{x}$ of size $n$.
+See also Table 2.
 3.1 The max-rank algorithm
@@
 In the conformal setting, $\mathbf{s}_{k}$ equates the set of nonconformity scores on $\mathcal{D}_{cal}$.
-We now desire finding adjusted significance levels $\alpha_{k}$ for all dimensions such that the FWER is globally controlled at target level $\alpha$.
+We now desire finding adjusted significance levels $\hat{\alpha}_{k}$ for all dimensions such that the FWER is globally controlled at target level $\alpha$.
 We algorithmically describe our proposed max-rank correction in Algo.
-2.
+2 (exemplary code is given in Listing 1).
 First, the empirical nulls of equal size are collected in a matrix of test statistics $\mathbf{S}=[s_{i,k}]\in\mathbb{R}^{n\times m}$, where $i\in[n]$ and $k\in[m]$.
@@
 By Eq.
-4, the search for a significance level $\alpha_{k}$ for each test is equivalent to computing the conformal quantile $\hat{Q}(1-\alpha;\mathbf{s}_{k})$ in the space of test statistics.
+4, testing at significance level $\alpha$ in each column is equivalent to computing the conformal quantile $\hat{Q}(1-\alpha;\mathbf{s}_{k})$ in the space of test statistics.
 This in turn equates determining the value $s_{t,k}$, where $t=\hat{Q}(1-\alpha;\mathbf{r}_{k})$ denotes a quantile in rank space ensuring the desired error control (Remark 1).
-max-rank adjusts these rank quantiles for FWER control by forming a composite empirical null $\mathbf{r}_{\max}$ via the $l^{\infty}$-norm across testing dimensions, and using the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ to inform globally corrected significance levels for each hypothesis test $k\in[m]$.
+The max-rank correction adjusts these rank quantiles for FWER control by forming a composite empirical null $\mathbf{r}_{\max}$ via the $l^{\infty}$-norm across tests, and using the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ to inform corrected significance levels $\hat{\alpha}_{k}$ for each hypothesis $k\in[m]$.
 Algorithm 2 Multiple testing correction via max-rank
 1:Input: Hypothesis tests $k=1,\dots,m$ and associated empirical nulls $\mathbf{s}_{1},\dots,\mathbf{s}_{m}$ of equal size $n$.
-2:Output: Adjusted quantiles with FWER control at level $\alpha$ as $(\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m})\in\mathbb{R}^{m}$.
+2:Output: Adjusted quantiles with FWER control at level $\alpha$ as $(\hat{Q}_{1-\hat{\alpha}_{1}},\dots,\hat{Q}_{1-\hat{\alpha}_{m}})\in\mathbb{R}^{m}$.
 3:Procedure:
@@
 Let $\mathbf{r}_{i}=(r_{i,1}\,\cdots\,r_{i,m})$ be the $i$-th row of $\mathbf{R}$, then $\mathbf{r}_{\max}=(\lVert\mathbf{r}_{1}\rVert_{\infty},\dots,\lVert\mathbf{r}_{n}\rVert_{\infty})^{T}=(\underset{1\leq k\leq m}{\max}|r_{1,k}|,\cdots,\underset{1\leq k\leq m}{\max}|r_{n,k}|)^{T}\in\mathbb{N}^{n}.$
-6:Compute the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ as the $\lceil(n+1)(1-\alpha)/n\rceil$-th empirical quantile of $\mathbf{r}_{\max}$.
+6:Compute the rank $r_{\max}=\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ as the $\lceil(n+1)(1-\alpha)\rceil/n$-th empirical quantile of $\mathbf{r}_{\max}$.
 7:Sort $\mathbf{S}$ column-wise ascending such that $s_{1,k}<\dots<s_{n,k}$ for every $k\in[m]$.
-8:Select the test statistic $s_{r_{\max},k}$ at rank $r_{\max}$ column-wise in $\mathbf{S}$ as the adjusted quantile $\hat{Q}_{1-\alpha,k}$ for every test dimension $k\in[m]$.
+8:Select the test statistic $s_{r_{\max},k}$ at rank $r_{\max}$ column-wise in $\mathbf{S}$ as the adjusted quantile $\hat{Q}_{1-\hat{\alpha}_{k}}$ for every test dimension $k\in[m]$.
 9:End Procedure
@@
 Our max-rank correction intuitively collapses multiple empirical nulls into a single ‘worst-case’ null for which the error-controlling quantile is computed.
-Thus, any potential dependencies across test statistics which are reflected in the rank ordering are efficiently aggregated in the ‘max-rank’ statistic $\mathbf{r}_{\max}$, and subsequently the selected rank quantile $r_{\max}$.
-By virtue of applying the $l^{\infty}$-norm, $r_{\max}$ ensures that adjusted significance levels provide FWER control.
+Thus, any potential dependencies across test statistics which are reflected in the rank ordering are efficiently aggregated in the ‘maximum rank’ statistic $\mathbf{r}_{\max}$, and subsequently the selected rank quantile $r_{\max}$.
+By virtue of applying the $l^{\infty}$-norm, $r_{\max}$ ensures that adjusted significance levels obey FWER control.
 If the empirical nulls are uncorrelated and the tested hypotheses can thus be considered independent, no benefit is obtained from max-rank’s aggregation step.
@@
 On the other hand, if the tests are perfectly correlated each empirical null exhibits identical rank ordering, and max-rank will provide the optimal quantile solution for each individual testing dimension.
-For any cases with positively dependent tests inbetween the two extremes, our approach ensures superior performance over Bonferroni, and scales well even as the number of hypotheses grows (see § 5).
+For any cases with positively dependent tests inbetween the two extremes, our approach ensures superior performance over Bonferroni, and scales well even as the number of hypotheses $m$ grows (see § 5).
+We highlight that no additional assumptions beyond those already underlying CP are required for the use of max-rank.
 3.2 Relation to Westfall & Young corrections
-A strong connection is apparent between max-rank and the max-T and min-P resampling-based corrections of Westfall and Young (1993), which offer similar FWER control.
+A strong connection exists between max-rank and the max-T and min-P resampling-based MHT corrections of Westfall and Young (1993), which offer similar FWER control.
 The max-T method provides adjusted p-values by comparing test statistics to the distribution of maximum statistics across hypotheses (Algo.
@@
 In both cases, existing dependency structure between tests is leveraged through repeated data resampling.
-Exploring these two corrections through the lens of our conformal prediction setting, we find that max-rank can be cast as a procedure that (1) naturally extends max-T to the rank space, and (2) describes a particular instance of min-P arising from our problem setting.
-As such, we argue that max-rank exhibits the advantages of both methods by (1) permitting its direct application to obtained test statistics as in max-T, whilst (2) remaining scale-invariant as in min-P.
+Exploring these two corrections through the lens of our conformal prediction setting, we find that max-rank can be cast as a procedure that (i) naturally extends max-T to the rank space, and (ii) describes a particular instance of min-P arising from our problem setting.
+As such, we argue that max-rank exhibits the advantages of both methods by permitting its direct application to obtained test statistics as in max-T, whilst maintaining scale invariance as in min-P.
 We next relate to each correction in more detail.
@@
 Our conformal setting directly provides empirical null distributions for every hypothesis by use of the calibration set.
-The computed nonconformity scores on $\mathcal{D}_{cal}$ leverage available labels to produce statistics satisfying the tested null $H_{0}:Y_{n+1}=y$, and are thus interpretable as an analog to the required “data resampling under the null” (Westfall and Young, 1993).
+The computed nonconformity scores on $\mathcal{D}_{cal}$ leverage available labels to produce statistics known to satisfy the tested null $H_{0}:Y_{n+1}=y$, and are thus interpretable as an analog to the required “data resampling under the null” (Westfall and Young, 1993).
+Similarly, CP’s exchangeability condition aligns with the exchangeability underlying the resampling step.
 It follows that the adjusted p-values in Algo.
@@
 Instead of adjusting the p-values as in Algo.
-3 we can directly adjust the significance level $\alpha$, which by Eq.
-4 translates to computing the quantile $\hat{Q}(1-\alpha;S_{\max})$ instead of $\hat{Q}(1-\alpha;S)$.
+3 we may also directly target an adjustment of the conformal quantile (Eq.
+4), which translates to computing $\hat{Q}(1-\alpha;S_{\max})$ instead of $\hat{Q}(1-\alpha;S)$.
 Comparing this to max-rank, we notice that the difference lies in its direct application to the maximum statistics, rather than translating them into rank space first (i.e., using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$).
@@
 By the same resampling interpretation as above, unadjusted p-values for each hypothesis in Algo.
-4 can be computed via conformal p-values in Eq.
-3.
+4 can be computed via conformal p-values (Eq.
+3).
 Using the rank definition in Eq.
-5, a given p-value $\hat{P}_{n+1}(y;S)$ can be expressed in terms of score ranking as $((n+1)-\text{rank}(s_{n+1};S))/(n+1)$ [^3] Observe that the permutation-based p-value in Eq.
-3 counts the fraction of values above the quantity of interest, whereas the rank counts values below it..
-Thus, computing the minimum p-values for min-P can be interpreted as taking the maximum over inverted and normalized ranks.
-Using Eq.
-4, we again translate adjusting the p-values as adjusting the significance level via $\hat{Q}(\alpha;\mathbf{p}_{\min})$ instead.
-Note that this quantile is now taken over (min.) p-values, and thus only permits testing in p-value space.
-In contrast, max-rank maintains the simplicity of operating directly on the statistics, facilitating the interpretation of testing decisions.
+5, a given p-value $\hat{P}_{n+1}(y;S)$ can be expressed in terms of score ranking as $((n+1)-\text{rank}(s_{n+1};S))/(n+1)$ [^2] Observe that the p-value in Eq.
+3 counts the fraction of values above $s_{n+1}$, whereas the rank counts values below it..
+Thus, computing the minimum p-values in min-P can be interpreted as the maximum over inverted and normalized ranks.
+Via Eq.
+4 we again translate adjusted p-values to a quantile adjustment using $\hat{Q}(\alpha;\mathbf{p}_{\min})$.
+Note that this quantile is now taken over p-values and only permits testing in p-value space.
+In contrast, max-rank keeps the simplicity of operating directly on statistics, facilitating the interpretation of testing decisions.
 4 THEORETICAL RESULTS
 We next briefly describe our theoretical results for max-rank, deferring details and proofs to Appendix A.
-The multiple testing problem is first rephrased in terms of multivariate copulas, and our two key results state that (1) max-rank provides a valid solution with FWER control, and (2) is assured to perform better than Bonferroni under positive dependency of tests.
+The multiple testing problem is first rephrased in terms of multivariate copulas, and our two key results state that (i) max-rank provides a valid MHT correction with FWER control, and (ii) is assured to perform better than Bonferroni under positive dependency.
 Recasting the multiple testing problem (§ A.1).
@@
 as $F_{\mathbf{X}}(\mathbf{x})$.
-Using the interpretation in Remark 1, determining adjusted significance levels $(\alpha_{1},\dots,\alpha_{m})$ with joint Type-I error control at level $\alpha$ is then the equivalent to finding a vector $\mathbf{x}_{*}\in\mathbb{R}^{m}$ satisfying $F_{\mathbf{X}}(\mathbf{x}_{*})\geq 1-\alpha$.
-By using Sklar’s Theorem (Sklar, 1973) to express the c.d.f.
+Using the interpretation in Remark 1, determining adjusted significance levels $(\hat{\alpha}_{1},\dots,\hat{\alpha}_{m})$ with joint Type-I error control at level $\alpha$ is then equivalent to finding a vector $\mathbf{x}_{*}\in\mathbb{R}^{m}$ satisfying $F_{\mathbf{X}}(\mathbf{x}_{*})\geq 1-\alpha$.
+By use of Sklar’s Theorem (Sklar, 1973) to express the c.d.f.
 in terms of a copula $C_{\mathbf{X}}$, and defining a vector of probabilities $(q_{*,1},\dots,q_{*,m})=(F_{X_{1}}(x_{*,1}),\dots,F_{X_{m}}(x_{*,m}))\in[0,1]^{m}$ restricted to a fixed value $q_{*}=q_{*,1}=\dots=q_{*,m}$, we can restate the problem as the search for a solution
@@
 6 the desired quantiles with FWER control are then computed as $\mathbf{x}_{*}=(F_{X_{1}}^{-1}(q_{*}),\dots,F_{X_{m}}^{-1}(q_{*}))\in\mathbb{R}^{m}$.
-These are equal to the quantiles $\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m}$ described in Algo.
+These equate the quantiles $\hat{Q}_{1-\hat{\alpha}_{1}},\dots,\hat{Q}_{1-\hat{\alpha}_{m}}$ described in Algo.
 2.
-Leveraging operations in the rank space, we show that max-rank provides such a solution $q_{\max}$.
+Leveraging operations in the rank space, we show that max-rank provides such a solution $q_{\max}$ to the problem.
 We state this in the following proposition:
@@
 Furthermore, the $\min$ implies that the associated quantiles $\mathbf{x}_{\max}$ are optimally tight in terms of coverage, under the condition that $q_{\max}=q_{\max,1}=\dots=q_{\max,m}$.
+[…]
+[…]
 Comparison to the Bonferroni correction (§ A.3).
@@
 6 ensures that we have $q_{\max}\leq q^{\perp}_{\max}\leq q_{\text{Bonf}}$ in the case of both independent and positively dependent test statistics.
-[…]
-[…]
 5 SIMULATION STUDY
@@
 We observe two key drawbacks for Bonferroni.
-Firstly, the correction becomes increasingly conservative as correlation between tests increases, with the adjusted level $\hat{\alpha}$ tending towards zero.
+Firstly, the correction becomes increasingly conservative as correlation between tests increases, with a FWER control level tending towards zero.
 Secondly, the rate accelerates as the number of tests increases.
-In contrast, max-rank suffers from neither of these pathologic behaviours and is able to tightly maintain FWER control at target level $\alpha$, ensuring good performance independently of both correlation strength and test dimensionality (Prop.
+In contrast, max-rank suffers from neither of these behaviours and tightly maintains FWER control at target level $\alpha$, ensuring consistent results independently of correlation strength and test dimensionality (Prop.
 1).
@@
 1 for $m=5$ that the FWER remains controlled.
-[…]
+FWER control and sample size (§ D.1).
+Since max-rank relates to Westfall and Young (1993) and is thus interpretable as a resampling-based correction, we anticipate the sample size $n$ (equivalent to the number of resamples) to affect FWER control, in particular when running many tests in parallel.
+Thus, we verify in Fig.
+4 adjusted significance levels for max-rank and Bonferroni across varying combinations of correlation $\rho$, sample size $n$ and test counts $m$.
+A direct correspondence between FWER control and the magnitudes of $m$ and $n$ is apparent: as we run more tests in parallel the need for additional samples increases in order to empirically maintain FWER control at target level $\alpha$ (e.g., roughly $n\geq 100\,m$ for $\rho=0$).
+Crucially, larger correlations suggest a lower need for additional samples as max-rank can exploit existing rank order information, whereas Bonferroni merely grows more conservative.
 6 APPLICATION TO CONFORMAL PREDICTION
-We now elaborate on how our proposed max-rank correction is useful for the conformal prediction setting introduced in § 2.1.
-First, we illustrate how multiple testing issues arise whenever multiple CP procedures are run in parallel.
-We then theoretically motivate why max-rank integrates well with the conformal setting, and experimentally validate its use in two conformal applications requiring such a correction.
+We further elaborate on the usefulness of our proposed max-rank correction for conformal prediction.
+First, we illustrate how multiple testing issues emerge when running multiple CP procedures in parallel.
+We then theoretically motivate why max-rank integrates well with the conformal design and its underlying conditions, and empirically validate its effectiveness in two CP applications that require a multiple testing correction.
 Multiple testing issues in conformal prediction.
-We previously motivated that running multiple conformal procedures at once equates permutation testing in parallel at significance level $\alpha$ (Eq.
-3).
+We previously discussed that running multiple conformal procedures at once equates permutation testing in parallel at a fixed significance level $\alpha$.
 By Eq.
-4 and Remark 1 this is the same as ensuring $(1-\alpha)$ coverage for an $m$-dimensional response $\mathbf{Y}_{n+1}=(Y_{n+1,1},\dots,Y_{n+1,m})$ simulatenously.
-Take $\mathbf{X}_{n+1}=(X_{n+1,1},\dots,X_{n+1,m})$ to be a matching multivariate input, then we directly observe how the coverage guarantee in Eq.
-2 is violated since
-$\mathbb{P}\left(\mathbf{Y}_{n+1}\in\hat{C}(\mathbf{X}_{n+1})\right)$ $=\mathbb{P}\left(\bigcap_{k=1}^{m}\left(Y_{n+1,k}\in\hat{C}(X_{n+1,k})\right)\right)$
-$\geq 1-\mathbb{P}\left(\bigcup_{k=1}^{m}\left(Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\right)\right)$
-$\geq 1-\sum_{k=1}^{m}\mathbb{P}\left(Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\right)$
-$=1-\sum_{k=1}^{m}\mathbb{P}\left(\hat{P}_{n+1,k}(Y_{n+1,k};S)\leq\alpha_{k}\right)$
-$\geq 1-\sum_{k=1}^{m}\alpha_{k}=1-m\alpha\quad\text{for }\alpha_{k}=\alpha\text{ fixed.}$. (7)
+4 and Remark 1 this is identical to ensuring a $(1-\alpha)$ coverage rate for each component of an $m$-dimensional response $\mathbf{Y}_{n+1}=(Y_{n+1,1},\dots,Y_{n+1,m})$.
+The same intuition is employed in our theoretical formulation for Eq.
+6.
+Take $\mathbf{X}_{n+1}=(X_{n+1,1},\dots,X_{n+1,m})$ to be a matching multivariate input, then we can directly observe how the coverage guarantee in Eq.
+2 can be violated since
+$\mathbb{P}\Big{(}$ $\mathbf{Y}_{n+1}\in\hat{C}(\mathbf{X}_{n+1})\Big{)}$
+$=\mathbb{P}\Big{(}\bigcap_{k=1}^{m}\big{(}Y_{n+1,k}\in\hat{C}(X_{n+1,k})\big{)}\Big{)}$
+$=1-\mathbb{P}\Big{(}\bigcup_{k=1}^{m}\big{(}Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\big{)}\Big{)}$
+$\geq 1-\sum_{k=1}^{m}\mathbb{P}\Big{(}Y_{n+1,k}\notin\hat{C}(X_{n+1,k})\Big{)}$
+$=1-\sum_{k=1}^{m}\mathbb{P}\Big{(}\hat{P}_{n+1,k}(Y_{n+1,k};S)\leq\alpha_{k}\Big{)}$
+$\geq 1-\sum_{k=1}^{m}\alpha_{k}=1-m\alpha,\quad\text{for }\alpha_{k}=\alpha\text{ fixed.}$. (7)
 That is, guaranteed coverage is limited to $1-m\alpha\leq 1-\alpha$ for any number of tests $m\in\mathbb{N}^{+}$, requiring a correction.
-A key insight is that due to their joint dependence on common calibration data $\mathcal{D}_{cal}$, any obtained conformal p-values are bound to exhibit some degree of positive dependency.
-This behaviour is referred to as positive regression dependent on a subset (PRDS) (Bates et al., 2022; Benjamini and Yekutieli, 2001; Chi et al., 2022), and highlights that any multiple testing correction relying on an independence assumption (such as Bonferroni) is bound to be overly conservative a priori.
+A key observation is that due to their joint dependence on common calibration data $\mathcal{D}_{cal}$, obtained conformal p-values in any single dimension are bound to exhibit some degree of positive dependency, a property known as positive regression dependent on a subset (PRDS) (Benjamini and Yekutieli, 2001; Bates et al., 2022; Chi et al., 2025).
+This highlights that any MHT correction relying on an independence assumption (such as Bonferroni) is bound to be overly conservative a priori, and our empirical results point towards a similar positive dependency across testing dimensions for CP settings.
 6.1 Using max-rank for conformal prediction
-Since we have intimately connected max-rank’s formulation to conformal prediction thus far, we may rightly assume that it integrates well with any CP algorithm following Algo.
+Since we have closely connected max-rank’s formulation to conformal prediction, one can rightly assume that it integrates well with any conformal procedure following Algo.
 1.
-In order to rigorously verify this fact, we assert that max-rank satisfies two key requirements imposed by conformal prediction: (1) the exchangeability condition on the underlying data is preserved, and (2) Eq.
+To rigorously assert this, we verify that max-rank meets two key requirements of conformal prediction: (i) the exchangeability condition on the underlying data is preserved, and (ii) Eq.
 7 is corrected in such a way that the conformal coverage guarantee from Eq.
 2 is upheld.
-We demonstrate in § A.4 that max-rank satisfies both conditions by looking at its two key operations, namely (1) operating in the rank space of nonconformity scores, and (2) applying a $\max$-operation to these ranks (or more generally, any set of exchangeable random variables).
-While the requirement on valid coverage directly follows from max-rank’s FWER-controlling property (Prop.
-1), an additional proof is given via rank coverage.
-We summarize these insights as follows:
+We demonstrate in § A.4 that max-rank satisfies both conditions by analysing its two key operations, namely (i) operating in the rank space of nonconformity scores, and (ii) applying a $\max$-operation on these ranks (or more generally, any set of exchangeable random variables).
+While the requirement on valid coverage technically directly follows from max-rank’s FWER-controlling property (Prop.
+1), an additional proof is provided via rank coverage.
+We summarize these insights in the following proposition:
 Proposition 3.
@@
 6.2 Conformal prediction experiments
-Finally, we showcase max-rank’s usefulness experimentally for two applications, conformal multi-target regression and conformal object detection.
-Experiment design.
-We consider applying conformal prediction to multi-target regression problems where $m$ real-valued and possibly dependent targets are predicted simultaneously.
-We aim to ensure that all $m$ targets are jointly covered with probability $(1-\alpha)$ by our conformal procedure, and consider a range of multiple testing corrections for that purpose: Bonferroni and its slightly more powerful Šidák-adjustment (Abdi et al., 2007), two copula-based corrections from Messoudi et al.
-(2021), max-T (Westfall and Young, 1993) and finally max-rank.
-We include the copula-based corrections because they also address settings of positive dependency, and since max-rank’s theoretical motivation via copulas provides an interesting link.
-For CP (Algo.
-1) we employ simple nonconformity scores of the form $|Y_{i,k}-\hat{f}(X_{i,k})|$ where $i\in[n],k\in[m]$, and construct the prediction intervals for test points as $\hat{C}(X_{n+1,k})=[\hat{f}(X_{n+1,k})-\hat{Q}_{1-\hat{\alpha},k},\,\hat{f}(X_{n+1,k})+\hat{Q}_{1-\hat{\alpha},k}]$, where $\hat{Q}_{1-\hat{\alpha},k}$ now denotes the conformal quantile at a possibly adjusted individual significance level $\hat{\alpha}_{k}$.
-We compare results with different corrections across two key conformal metrics (Vovk et al., 2005): empirical coverage as a measure of validity to verify Eq.
-2, and average prediction interval width to assess the efficiency of the solution.
-See also Appendix D.
 Figure 3: Empirical coverage (left) and mean prediction interval width (right) for different corrections ($\alpha=0.1$) on BDD100k.
@@
 deviation across $100$ random trials.
-max-rank consistently provides strong results (best results, second best).
+max-rank consistently provides strong results (best, second best results).
+Finally, we showcase max-rank’s usefulness experimentally for two applications, conformal multi-target regression and conformal object detection.
+Our code is available at https://github.com/alextimans/max-rank.
+Experiment design.
+We consider applying conformal prediction to multi-target regression problems where $m$ real-valued and possibly dependent targets are predicted simultaneously.
+We aim to ensure that all $m$ targets are jointly covered with probability $(1-\alpha)$ by our conformal procedure, and consider a range of multiple testing corrections for that purpose: Bonferroni and its slightly more powerful Šidák-adjustment (Abdi et al., 2007), two copula-based corrections from Messoudi et al.
+(2021), max-T (Westfall and Young, 1993) and finally max-rank.
+We include the copula-based corrections because they also address settings of positive dependency, and since max-rank’s theoretical motivation via copulas provides an interesting link.
+For CP (Algo.
+1) we employ nonconformity scores of the simple residual form $s_{i,k}=|Y_{i,k}-\hat{f}(X_{i,k})|$ where $i\in[n],k\in[m]$, and construct the prediction intervals for test points as $\hat{C}(X_{n+1,k})=[\hat{f}(X_{n+1,k})-\hat{Q}_{1-\hat{\alpha}_{k}},\,\hat{f}(X_{n+1,k})+\hat{Q}_{1-\hat{\alpha}_{k}}]$, where $\hat{Q}_{1-\hat{\alpha}_{k}}$ again denotes the conformal quantile at adjusted individual significance level $\hat{\alpha}_{k}$ using a MHT correction.
+We compare results with different corrections across two key conformal metrics (Vovk et al., 2005): empirical coverage as a measure of validity to verify Eq.
+2, and average prediction interval width to assess the efficiency of the solution.
+See also Appendix D.
 Conformal multi-target regression.
@@
 Our results for the usual target coverage of $(1-\alpha)=90\%$ are displayed in Table 1.
-Firstly, we observe that all corrections provide valid coverage, whereas without we (strongly) undercover.
+Firstly, we observe that all corrections provide valid coverage, whereas without we strongly undercover.
 Among corrections, Bonferroni and Bonferroni-Šidák provide overly conservative results, and copula methods perform generally well.
@@
 7 DISCUSSION
-We introduced max-rank, an efficient MHT correction for testing under positive dependency which tailors Westfall and Young (1993)’s corrections for conformal prediction.
-While we provide strong theoretical and empirical results, our method’s practicality is restricted to (1) data dependencies that become apparent through rank ordering, (2) settings of positive dependency, since we do not benefit from rank information under negative dependency and thus do not improve upon Bonferroni, and (3) its permutation-based, data-driven nature necessitating a sufficient ratio of test statistics $n$ to test dimensions $m$ for empirical FWER control.
+We introduce max-rank, an efficient MHT correction for testing under positive dependency which adapts Westfall and Young (1993)’s corrections to conformal prediction.
+While we provide strong theoretical and empirical results, our method’s benefits are mostly restricted to (i) data dependencies that become apparent through rank ordering, (ii) settings of positive dependency, since we do not benefit from rank information under negative dependency and thus do not improve upon Bonferroni, and (iii) its permutation-based, data-driven nature necessitates a sufficient ratio of test statistics $n$ to test dimensions $m$ for empirical FWER control.
 Future work includes establishing stronger theoretical results, such as on convergence rates to target level $\alpha$, the dependence between quantities $m$, $n$, and $\alpha$, and its relationship to the empirical copula.
-Of course, we may also experimentally extend to further applications such as multi-step time series (Sun and Yu, 2023).
-Our work contributes to the ongoing research on multiple testing under dependency, all the while addressing existing needs in predictive settings using conformal prediction.
+Naturally, we may also experimentally extend to further applications such as multi-step time series (Sun and Yu, 2023).
+We hope that max-rank contributes to the ongoing research on multiple testing under dependency, all the while addressing practical needs in predictive settings using conformal prediction.
 Acknowledgements
-We thank Aaditya Ramdas and Jelle Goeman for making us aware of connections to the Westfall & Young testing procedures.
+We thank Aaditya Ramdas and Jelle Goeman for making us aware of connections to the Westfall & Young testing procedures, and Drew Prinster for helpful feedback on an earlier draft.
 This project was generously supported by the Bosch Center for Artificial Intelligence.
@@
 Christoph-Nikolas Straehle developed the max-rank algorithm and proved several theoretical properties.
-Kaspar Sakmann, Christian Naesseth and Eric Nalisnick provided project guidance as well as technical and writing feedback.
+Kaspar Sakmann, Christian A.
+Naesseth and Eric Nalisnick provided project guidance and helpful feedback.
+CHECKLIST
+  1. 1.
+    For all models and algorithms presented, check if you include:
+    1. (a)
+      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes] Algorithmic descriptions are included in the paper, the full proofs are included in the appendix.
+    2. (b)
+      An analysis of the properties and complexity (time, space, sample size) of any algorithm.
+[Yes] We discuss the algorithmic properties and empirically verify sample size and runtime requirements, but do not include a formal theoretical analysis.
+    3. (c)
+      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes] Our code is publicly available at https://github.com/alextimans/max-rank.
+  2. 2.
+    For any theoretical claim, check if you include:
+    1. (a)
+      Statements of the full set of assumptions of all theoretical results. [Yes] The full proofs are included in the appendix.
+    2. (b)
+      Complete proofs of all theoretical results. [Yes] The full proofs are included in the appendix.
+    3. (c)
+      Clear explanations of any assumptions. [Yes] The full proofs are included in the appendix.
+  3. 3.
+    For all figures and tables that present empirical results, check if you include:
+    1. (a)
+      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL).
+[Yes] We provide experimental design details in the paper and the appendix, and the code is publicly available.
+    2. (b)
+      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
+    3. (c)
+      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
+    4. (d)
+      A description of the computing infrastructure used.
+(e.g., type of GPUs, internal cluster, or cloud provider).
+[Yes] All results are reproducible on a single consumer-grade GPU (e.g., NVIDIA RTX 4090).
+max-rank is computationally light-weight and is easily runnable on CPU for smaller experiments.
+  4. 4.
+    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
+    1. (a)
+      Citations of the creator If your work uses existing assets. [Yes]
+    2. (b)
+      The license information of the assets, if applicable. [Yes] Our code is publicly released under an AGPL-3.0 license.
+    3. (c)
+      New assets either in the supplemental material or as a URL, if applicable. [Yes] The code is available via an URL.
+    4. (d)
+      Information about consent from data providers/curators. [Not Applicable] Used datasets are synthetic or publicly available.
+    5. (e)
+      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
+  5. 5.
+    If you used crowdsourcing or conducted research with human subjects, check if you include:
+    1. (a)
+      The full text of instructions given to participants and screenshots. [Not Applicable]
+    2. (b)
+      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
+    3. (c)
+      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
+Max-Rank: Efficient Multiple Testing for Conformal Prediction
+– Supplementary Materials –
+The appendix is organized as follows:
+  • Mathematical details and proofs in Appendix A
+  • Algorithmic details on max-T and min-P in Appendix B
+  • Additional background on FDR control and the Bonferroni correction in Appendix C
+  • Additional experiment details for our simulation study, multiple testing baselines, metrics, and conformal experiments in Appendix D
+Notation Type Description
+$n$ $\in\mathbb{N}^{+}$ Nr. of calibration samples, nr. of test statistics, index $i=1,\dots,n$
+$m$ $\in\mathbb{N}^{+}$ Nr. of conformal procedures, nr. of test dimensions, index $k=1,\dots,m$
+$\alpha,\alpha_{k}$ $\in(0,1)$ Target miscoverage rate, target significance level (global, in the $k$-th dimension)
+$\hat{\alpha},\hat{\alpha}_{k}$ $\in(0,1)$ Adjusted or corrected significance level (global, in the $k$-th dimension)
+$\hat{f}(\cdot)$ $\mathcal{X}\rightarrow\mathcal{Y}$ Pre-trained prediction model
+$s(\cdot)$ $\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ Conformal scoring function
+$s_{i},s_{i,k}$ $\in\mathbb{R}$ Nonconformity score for the $i$-th sample, test statistic (in the $k$-th dimension)
+$r_{i},r_{i,k}$ $\in\mathbb{N}$ Rank of nonconformity score for the $i$-th sample (in the $k$-th dimension)
+$S$ $\subset\mathbb{R}$ Set of scores $\{s_{1},\dots,s_{n}\}$ given a single dimension (equiv. to fixed $\mathbf{s}_{k}$ ignoring ordering)
+$\hat{C}(X_{n+1})$ $\subseteq\mathcal{Y}$ Conformal prediction set for test input $X_{n+1}$
+$\hat{P}_{n+1}(y;S)$ $\in[0,1]$ Conformal p-value for the null hypothesis $H_{0}:Y_{n+1}=y$ based on $S$
+$\hat{Q}(1-\alpha;S),\,\hat{Q}_{1-\alpha}$ $\in\mathbb{R}$ Conformal quantile at target coverage $(1-\alpha)$ based on $S$
+$\hat{Q}(1-\alpha_{k};\mathbf{s}_{k}),\hat{Q}_{1-\alpha_{k}}$ $\in\mathbb{R}$ Conformal quantile at target coverage $(1-\alpha_{k})$ based on $\mathbf{s}_{k}$ (i.e., in the $k$-th dimension)
+$\mathbf{s}_{k}$ $\in\mathbb{R}^{n}$ Vector of scores, test statistics, empirical null distribution (in the $k$-th dimension)
+$\mathbf{r}_{k}$ $\in\mathbb{N}^{n}$ Vector of score ranks, empirical null distribution in rank space (in the $k$-th dimension)
+$\mathbf{S}$ $\in\mathbb{R}^{n\times m}$ Matrix of test statistics with columns $\mathbf{s}_{1},\dots,\mathbf{s}_{m}$
+$\mathbf{R}$ $\in\mathbb{N}^{n\times m}$ Matrix of column-wise ranks of $\mathbf{S}$
+$\mathbf{r}_{\max}$ $\in\mathbb{N}^{n}$ Vector of max. score ranks, composite empirical null distribution in rank space
+$r_{\max}$ $\in\mathbb{N}$ The rank value $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$ representing the corrected significance levels $\hat{\alpha},\hat{\alpha}_{k}$
+Table 2: Notation table with clarifications for key quantities presented in the main paper.
 Appendix A MATHEMATICAL DETAILS
-This section includes the details of our theoretical results for max-rank (Prop. 1 and Prop. 2 in § 4), as well as its use for conformal prediction (Prop. 3 in § 6.1).
+This section includes notation clarification in Table 2, and the details of our theoretical results for max-rank (Prop.
+1 and Prop.
+2 in § 4), as well as its use for conformal prediction (Prop.
+3 in § 6.1).
 A.1 Recasting the multiple testing problem
@@
 Note that now the elements of $\mathbf{x_{*}}$ may differ from each other, since the marginals are not necessarily identical.
-The obtained quantiles in $\mathbf{x_{*}}$ are equal to the adjusted quantiles $\hat{Q}_{1-\alpha,1},\dots,\hat{Q}_{1-\alpha,m}$ described in Algo.
+The obtained quantiles in $\mathbf{x_{*}}$ are equal to the adjusted quantiles $\hat{Q}_{1-\hat{\alpha}_{1}},\dots,\hat{Q}_{1-\hat{\alpha}_{m}}$ described in Algo.
 2.
@@
 2 to our problem setting in § A.1.
-Consider the matrix of test statistics $\mathbf{S}\in\mathbb{R}^{n\times m}$ first introduced in § 3, and denote by $S_{k}$ its $k$-th column corresponding to the $k$-th empirical null.
-We then associate $S_{k}$ with the $k$-th dimension of $\mathbf{X}$, i.e., the elements $s_{i,k}\in S_{k},\,i\in[n]$ match to $n$ realizations $\{x_{1,k},\dots,x_{n,k}\}$ of $X_{k}$, and $S_{k}$ thus provides an empirical estimate of the marginal $F_{X_{k}}$.
+Consider the matrix of test statistics $\mathbf{S}\in\mathbb{R}^{n\times m}$ first introduced in § 3, and denote by $S_{k}$ its $k$-th column corresponding to the $k$-th empirical null (in set notation, rather than as a vector $\mathbf{s}_{k}$).
+We then associate $S_{k}$ with the $k$-th dimension of $\mathbf{X}$, i.e., the elements $s_{i,k}\in S_{k},\,i\in[n]$ match with the set of $n$ realizations $\{x_{1,k},\dots,x_{n,k}\}$ of $X_{k}$, and $S_{k}$ thus provides an empirical estimate of the marginal $F_{X_{k}}$.
 Next, we evaluate the element-wise ranks of $\mathbf{x_{*}}$ under $\mathbf{S}$.
@@
 Intuitively, ensuring that the condition holds jointly for ranks in all dimensions is equivalent to satisfying it in the ‘worst’ one.
-$F_{\max\mathbf{\tilde{R}}}$ is precisely the space in which max-rank operates after the $l^{\infty}$-norm is applied, and in which following Algo.
+$F_{\max\mathbf{\tilde{R}}}$ is precisely the space in which max-rank operates after the $l^{\infty}$-norm is applied (matching the composite null $\mathbf{r}_{\max}$ in the paper), and in which following Algo.
 2 a rank $\tilde{r}_{\max}=r_{\max}/n=q_{\max}$ with probability mass $(1-\alpha)$ below it is computed.
@@
 Note that if the restriction is lifted, marginally tighter individual quantiles $x_{\max,k}$ are potentially obtainable in some dimensions.
-However, it becomes unclear how the required algorithmic modifications to max-rank would continue to uphold FWER control.
+However, it becomes unclear how the required algorithmic modifications to max-rank would continue to formally uphold FWER control.
 We summarize our arguments by restating our proposition from § 4, which follows from the above:
@@
 A.4 Using max-rank for conformal prediction
-Two key conditions are imposed on any multiple testing correction that integrates with conformal prediction: (1) the exchangeability condition on the underlying data is preserved, and (2) Eq.
+Two key conditions are imposed on any multiple testing correction that integrates with conformal prediction: (i) the exchangeability condition on the underlying data is preserved, and (ii) Eq.
 7 is corrected in such a way that the conformal coverage guarantee is upheld.
-We next show that max-rank satisfies both conditions by looking at its two key operations, namely (1) operating in the rank space of nonconformity scores, and (2) applying a $\max$-operation to these ranks.
+We next show that max-rank satisfies both conditions by looking at its two key operations, namely (i) operating in the rank space of nonconformity scores, and (ii) applying a $\max$-operation to these ranks.
 We first address the exchangeability requirement, followed by validity.
@@
 2 is maintained also when operating in the rank space, we employ the following corollary:
-Corollary 1 ((Kuchibhotla, 2020), Corr. 1).
+Corollary 1 (Kuchibhotla (2020), Cor. 1).
 Under assumptions of Thm. 1 we have that
@@
 2 we have for any $k\in[m]$ that $\hat{Q}(1-\alpha;\mathbf{r}_{\max})\geq\hat{Q}(1-\alpha;\mathbf{r}_{k})$ holds.
-Clearly, a prediction set $\hat{C}_{k}$ constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{k})$ is then a subset of $\hat{C}_{\max}$ constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$, i.e., $\hat{C}^{\max}\supseteq\hat{C}^{k}$.
+Clearly, a prediction set $\hat{C}_{k}$ for the $k$-th dimension constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{k})$ is then a subset of the prediction set $\hat{C}_{\max}$ in the $k$-th dimension constructed using $\hat{Q}(1-\alpha;\mathbf{r}_{\max})$, i.e., $\hat{C}_{\max}\supseteq\hat{C}_{k}$.
 Since by Eq.
@@
 Appendix B ALGORITHMIC DETAILS
+The algorithmic descriptions for Westfall and Young (1993)’s max-T and min-P multiple testing corrections are given in Algo.
+3 and Algo.
+4 respectively.
+The main paper contains the algorithms for split conformal prediction in the general case (Algo.
+1) and our correction max-rank (Algo.
+2).
+A simple implementation of max-rank takes only a few lines of code, and we provide an example in Listing 1.
+In this example we simulate perfect correlation ($\rho=1.0$) and observe how max-rank exploits the rank ordering and therefore does not adjust the initial significance level (i.e., $\hat{\alpha}=\alpha=0.1$) which returns the (identical) optimal quantile in each test dimension ($m=5$).
+In contrast, Bonferroni would adjust to $\hat{\alpha}_{\text{Bonf}}=\alpha/m=0.02$ in each dimension.
+For $\rho\approx 0$ the same quantiles as Bonferroni are recovered, whereas $\rho\in(0,1)$ provides tighter quantiles (with varying degree).
 Algorithm 3 Multiple testing correction via max-T (Westfall and Young (1993))
 1:Input: Hypothesis tests $k=1,\dots,m$ and associated test statistics $(T_{1},\dots,T_{m})\in\mathbb{R}^{m}$, a resampling counter $i\in[n]$, and empty vector of resampled test statistics $\mathbf{t}_{\max}\in\mathbb{R}^{n}$.
-2:Output: Adjusted p-values $(P^{\text{adj}}_{1},\dots,P^{\text{adj}}_{m})\in\mathbb{R}^{m}$ with FWER control at level $\alpha$.
+2:Output: Adjusted p-values $(P^{\text{adj}}_{1},\dots,P^{\text{adj}}_{m})\in[0,1]^{m}$ with FWER control at level $\alpha$.
 3:Procedure:
@@
 Recompute test statistics for $i$-th resampling and store maximum statistic: $\mathbf{t}_{\max,i}=\max_{1\leq k\leq m}T_{i,k}$.
-5:Compute adjusted p-value by comparing observed statistics to the maximum statistics under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\mathbf{t}_{\max,i}\geq T_{k}]$.
+5:Compute adjusted p-value by comparing observed statistics to the maximum statistics under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}[T_{k}\leq\mathbf{t}_{\max,i}]$.
 6:Decision rule: $\forall k\in[m]$, we cannot reject the null if $P^{\text{adj}}_{k}>\alpha$.
@@
 5:For $i=1,\dots,n$: Resample data, e.g., via permutation or bootstrapping. Recompute p-values for $i$-th resampling and store the minimum p-value: $\mathbf{p}_{\min,i}=\min_{1\leq k\leq m}P_{i,k}$.
-6:Compute adjusted p-value by comparing observed p-values to the minimum p-values under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\mathbf{p}_{\min,i}\leq P_{k}]$.
+6:Compute adjusted p-value by comparing observed p-values to the minimum p-values under the null: $P^{\text{adj}}_{k}=\frac{1}{n}\sum_{i=1}^{n}\mathds{1}[P_{k}\geq\mathbf{p}_{\min,i}]$.
 7:Decision rule: $\forall k\in[m]$, we cannot reject the null if $P^{\text{adj}}_{k}>\alpha$.
 8:End Procedure
-The algorithmic descriptions for Westfall and Young (1993)’s max-T and min-P multiple testing corrections are given in Algo.
-3 and Algo.
-4 respectively.
-The main paper contains the algorithms for split conformal prediction in the general case (Algo.
-1) and our correction max-rank (Algo.
-2).
+Listing 1: Exemplary Python code for the max-rank correction.
+import numpy as np
+alpha = 0.1; n = 1000; m = 5; corr = 1.0 # params
+# simulate correlated score matrix
+cov = np.full((m, m), corr)
+np.fill_diagonal(cov, 1)
+S = np.random.multivariate_normal(np.zeros(m), cov, n)
+# max-rank correction
+R = np.argsort(np.argsort(S, axis=0), axis=0) # column-wise rank matrix
+r_max = np.max(R, axis=1) # row-wise max-rank vec(r)_max
+q = np.ceil((1 - alpha) * (n + 1)) / n # conformal target coverage level
+rank_q = np.quantile(r_max, q, axis=0, method="higher") # quantile r_max
+adj_quantile = np.sort(S, axis=0)[rank_q] # final corrected quantiles
+print(f"Adjusted significance level: {(1 - rank_q / n):.2f}") # 0.10
+print(f"Adjusted quantiles: {adj_quantile}") # e.g. [1.32, 1.32, 1.32, 1.32, 1.32]
 Appendix C ADDITIONAL BACKGROUND
+We provide additional details on FDR control (§ C.1) and the Bonferroni correction (§ C.2).
 C.1 FDR control
@@
 In contrast to classical corrections with FWER control which assume independence, FDR-controlling procedures are generally more robust to dependencies between tests.
+Beyond well-known FWER and FDR control, extensions such as $k$-FWER (Romano and Wolf, 2007; Janson and Su, 2016) – which bounds the probability of committing more than $k$ Type-I errors (recovering FWER for $k=1$) – and false discovery exceedance (Genovese and Wasserman, 2004) – which controls tail probabilities of the false discovery proportion (recovering FDR for the expectation) – have also been explored.
 C.2 FWER control via the Bonferroni correction
@@
 Appendix D ADDITIONAL EXPERIMENT DETAILS
+We give additional experiment details for our simulation study (§ D.1), multiple testing baselines (§ D.2), metrics (§ D.3), and conformal experiments (§ D.4, § D.5).
 D.1 Simulation study: FWER control and sample size
-Figure 4: We examine global adjusted significance levels $\hat{\alpha}$ against the desired FWER control level $\alpha=0.05$ for varying correlations $\rho$, number of tests $m$ and sample size $n$.
+Figure 4: We examine global adjusted significance levels $\hat{\alpha}$ against the desired FWER control level $\alpha=0.05$ for varying correlations $\rho$, number of tests $m$ and sample size $n$ (shading denotes std.
+deviation across 10 random trials).
 Empirical FWER control rates are directly related to the magnitudes of $m$ and $n$, suggesting that larger number of parallel tests require larger sample sizes.
 We also see that Bonferroni is consistently more conservative than max-rank across all combinations.
-[…]
-[…]
-The observed interactions between quantities seem consistent and suggest a meaningful relationship that warrants future theoretical investigation, as mentioned in § 7.
+Since max-rank connects to Westfall and Young (1993)’s corrections as a resampling-based method, we anticipate that its data-driven nature requires an adequate ratio of tests counts $m$ to sample sizes $n$ per test (equivalent to the number of resamples or permutations) in order to empirically provide FWER control.
+Thus, we verify in Fig.
+4 adjusted significance levels for max-rank and Bonferroni across varying combinations of correlation $\rho$, test counts $m$, and samples $n$.
+Specifically, we consider $\rho\in\{0,0.25,0.5,0.75,1.0\}$, $m\in[1,1000]$, and $n\in\{100,500,1000\}$.
+In particular, this also includes high-dimensional settings where $m>n$.
+Our results confirm the anticipated dependencies.
+Specifically, empirical FWER control is not guaranteed when the ratio of samples to test dimensions is too low, including the settings where $m>n$.
+These violations affect both max-rank and Bonferroni, albeit our correction tends to violate sooner since it operates closer to target level.
+However, FWER control rates improve substantially as both sample size and correlation increases.
+Intuitively, higher correlation effectively reduces the amount of independent rank order information across tests that needs to be accounted for.
+Overall, we observe that empirical FWER control is approximately maintained when $n\geq 100\cdot m$ in the least favorable setting ($\rho=0$).
+However, there are cases where this rate can be exceeded (e.g., $n=100$ and $m=5$ holds) or fall short (e.g., $n=1000$ and $m=10$ undercorrects slightly).
+Finally, throughout all considered combinations Bonferroni is always more conservative than max-rank (in line with Prop.
+2), and its disadvantage to max-rank increases substantially as the correlation goes up.
+The observed sample requirements seem reasonable given that a reliable empirical null distribution must be established for each test dimension, and are somewhat in the range of calibration set sizes commonly used in conformal prediction.
+For instance, Angelopoulos et al.
+(2023) recommend at least 1000 calibration samples for a robust conformal procedure even in a single dimension, while Messoudi et al.
+(2021) (see Table 2 in their paper) consider varying sample requirements for multi-target regression such as $(n,m)\in\{(2100,3),(24160,4),(910,8)\}$.
+The impact of overly small sample sizes on results is evident both in their work (e.g., Fig.
+3 in their paper) and ours (Table 1, for the scpf dataset).
+In general, the observed interactions between quantities seem consistent and suggest a meaningful relationship that warrants future theoretical investigation, as mentioned in § 7.
 It would be desirable to quantify the expected FWER control achievable for a given significance level $\alpha$, sample size $n$ and number of tests $m$ in analytical form.
@@
 Empirical Copula. The empirical copula is the copula-equivalent of an empirical, non-parametric estimator of the joint cumulative distribution. It evaluates to
-$C_{\text{Emp}}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\prod_{k=1}^{m}\mathbbm{1}[u_{i,k}\leq u_{k}],$ (21)
-where $\mathbbm{1}[\cdot]$ is the indicator function, $u_{k}$ is a component of $\mathbf{u}$ as above (equivalent to $1-\alpha$), and $u_{i,k}$ is the cumulative probability of the $i$-th sample under its marginal c.d.f.
+$C_{\text{Emp}}(\mathbf{u})=\frac{1}{n}\sum_{i=1}^{n}\prod_{k=1}^{m}\mathds{1}[u_{i,k}\leq u_{k}],$ (21)
+where $\mathds{1}[\cdot]$ is the indicator function, $u_{k}$ is a component of $\mathbf{u}$ as above (equivalent to $1-\alpha$), and $u_{i,k}$ is the cumulative probability of the $i$-th sample under its marginal c.d.f.
 in the $k$-th dimension (estimated empirically), i.e., $u_{i,k}=\hat{F}_{X_{k}}(x_{i,k})$.
-We can think of the copula as empirically estimating the fraction of samples jointly satisfying the condition $\mathbbm{1}[u_{i,k}\leq u_{k}]$ in all dimensions, since only in such cases does the product evaluate to one.
-As there is no analytical solution (Messoudi et al., 2021) we have to search through a grid of possible adjusted significance levels $\hat{\alpha}\in(0,1)$, evaluate them under the copula, and select the largest one satisfying target coverage $C_{\text{Emp}}(1-\hat{\alpha},\dots,1-\hat{\alpha})\geq(1-\alpha)$, thus directly addressing the problem in Eq.
+We can think of the copula as empirically estimating the fraction of samples jointly satisfying the condition $\mathds{1}[u_{i,k}\leq u_{k}]$ in all dimensions, since only in such cases does the product evaluate to one.
+As there is no analytical solution (Messoudi et al., 2021) we have to search through a grid of possible adjusted (global) significance levels $\hat{\alpha}\in(0,1)$, evaluate them under the copula, and select the largest one satisfying target coverage $C_{\text{Emp}}(1-\hat{\alpha},\dots,1-\hat{\alpha})\geq(1-\alpha)$, thus directly addressing the problem in Eq.
 6 (note that $\min q_{*}\Leftrightarrow\max\hat{\alpha}$).
@@
 Thus the average empirical coverage is given by
-$\text{Coverage}=\frac{1}{n_{t}\,m}\sum_{j=n+1}^{n+n_{t}}\mathbbm{1}\left[\bigcap_{k=1}^{m}(Y_{j,k}\in\hat{C}(X_{j,k}))\right],$ (22)
-where $\mathbbm{1}[\cdot]$ is the indicator function.
+$\text{Coverage}=\frac{1}{n_{t}\,m}\sum_{j=n+1}^{n+n_{t}}\mathds{1}\left[\bigcap_{k=1}^{m}(Y_{j,k}\in\hat{C}(X_{j,k}))\right],$ (22)
+where $\mathds{1}[\cdot]$ is the indicator function.
 We note that the empirical coverage is a random quantity parametrized by a coverage distribution, and can slightly deviate from target coverage $(1-\alpha)$ based on factors such as the size of the calibration data $|\mathcal{D}_{cal}|$ (Vovk et al., 2005).
@@
 Results are averaged across 100 random trials.
-max-rank consistently provides strong results also on a per-class basis (best results, second best).
+max-rank consistently provides strong results also on a per-class basis (best, second best results).
 We have varying object counts per class, equivalent to the number of per-class calibration samples $n$.
-By class from top to bottom: 44’960 (avg.), 35’748, 2032, 280’277, 883, 5151, 10’975, and 21’950.
+From top to bottom by class: 44’960 (Class avg.), 35’748, 2032, 280’277, 883, 5151, 10’975, and 21’950.
 Figure 5: A visual example of the constructed conformal prediction intervals with max-rank on a test image for the class ‘car’ (and not ‘truck’).